### v1

# Direct Superpoints Matching for Fast and Robust Point Cloud Registration

Abstract
Although deep neural networks endow the downsampled superpoints with discriminative feature representations, directly matching them is usually not used alone in state-of-the-art methods, mainly for two reasons. First, the correspondences are inevitably noisy, so RANSAC-like refinement is usually adopted. Such ad hoc postprocessing, however, is slow and not differentiable, which can not be jointly optimized with feature learning. Second, superpoints are sparse and thus more RANSAC iterations are needed. Existing approaches use the coarse-to-fine strategy to propagate the superpoints correspondences to the point level, which are not discriminative enough and further necessitates the postprocessing refinement. In this paper, we present a simple yet effective approach to extract correspondences by directly matching superpoints using a global softmax layer in an end-to-end manner, which are used to determine the rigid transformation between the source and target point cloud. Compared with methods that directly predict corresponding points, by leveraging the rich information from the superpoints matchings, we can obtain more accurate estimation of the transformation and effectively filter out outliers without any postprocessing refinement. As a result, our approach is not only fast, but also achieves state-of-the-art results on the challenging ModelNet and 3DMatch benchmarks. Our code and model weights will be publicly released.
1 Introduction
Point cloud registration refers to the task of aligning two partially overlapping point clouds into a shared coordinate system. In this paper, we tackle the problem of rigid registration where the goal is to determine the optimal transformation matrix, including rotation and translation, from one point cloud (source) to the other (target). It has attracted a lot of research interest due to its broad applications in SLAM (Simultaneous Localization and Mapping) [16, 50], autonomous driving [34, 37], 3D reconstruction [26, 22], etc.
Figure 1: Our approach focuses on directly matching the predicted superpoints by computing correlation between them. The bold dots show the superpoints with highest correlation which are used as correspondences to correctly register the point cloud pair.
A prevailing paradigm to solve the registration task is to leverage the correspondences of superpoints across two point clouds, which can be obtained either using separate keypoint detectors [1, 4, 13, 24] to capture salient and distinctive points or regions within a point cloud. With the rapid advancement in representation learning in point clouds using deep neural networks [42, 43, 51], the keypoint-free approaches have gained significant attention, where the point cloud is downsampled into superpoints. Although they may not be designed to capture the salient keypoints in the point cloud, the learned representations endow the superpoints with discriminative power so they can be matched across the source and target point clouds. The transformation matrix can then be obtained based on the superpoints correspondences using the Kabsh-Umeyama Algorithm [53, 28].
Direct superpoint matching is typically not used by itself in existing approaches for point cloud registration for two reasons. First, the correspondences of superpoints inevitably contain errors, either because the point clouds contain noisy sensory data or because some of the correspondences are simply incorrect. As a result, postprocessing refinement is usually needed, for instance, using RANSAC to prune out the outliers [13, 14, 68]. However, RANSAC is inherently slow and not differentiable, and thus cannot be integrated into the training step. Second, superpoints are essentially sparse, particularly because of downsampling in neural networks. Therefore, more RANSAC iterations are needed as the number of inliers (i.e., correct correspondences) is small, which imposes a significant computation burden. To this end, the coarse-to-fine registration scheme is usually adopted in state-of-the-art approaches [44, 66, 38], where superpoints correspondences serve as coarse correspondences only. They are then propagated to the point level, forming finer-level correspondences. Since point-wise features only capture local information, they lack enough discriminative power. Consequently, the noise in the point-wise correspondences is more prominent, which further necessitates the additional refinement step. Although the Local to Global Registration (LGR) module introduced in [44] improves the speed, which works similarly to RANSAC in an iterative manner, the selection and refinement of the transformation still remain non-differentiable. The entire model thus can not be optimized jointly, leading to inferior accuracy.
In this paper, we present a simple yet effective approach for rigid point cloud registration by directly matching the superpoints. Specifically, superpoints features are first obtained from the KPConv backbone [51], which are then enhanced using Transformer blocks with interleaved multi-head self and cross attention modules [64], enabling effective learning of the structural information of from the target to the source point cloud. The enhanced superpoint features are finally matched using a Global Softmax layer inspired by the 2D matching work [60], which generates a correlation matrix capturing the pairwise relationships between the keypoints from the source and target point clouds. But unlike [60], we fully harness the rich information in the correlation matrix. We employ a differentiable variant of the Kabsh-Umeyama Algorithm [53, 28], considering the strengths of matchings, where the weights of the SVD (Singular Value Decomposition) are the correlation scores between two corresponding superpoints. By examining the correlation scores of the superpoints, we can easily identify the highly reliable correspondences without resorting to the time-consuming RANSAC or the non-differentiable LGR module [44]. Our approach is fully differentiable and can be trained in an end-to-end manner, which enables joint optimization of the feature representation learning, superpoints matching, and transformation estimation, leading to improved alignment results. Moreover, our approach does not require ad-hoc postprocessing, e.g., RANSAC, it runs fast, making it potentially more useful in practice. Compared with the recent end-to-end method [64], our approach directly matches the superpoints. With the rich information obtained from the matching step, our approach can get a more accurate estimation of the transformation between the point clouds robustly by filtering out the outliers effectively.
We run experiments on standard benchmarks, including ModelNet [58] and 3DMatch [68], and achieve state-of-the-art results. Extensive ablation studies validate the effectiveness of each module in our proposed approach. The key contributions of the paper can be summarized as follows:
  • Efficiency. Our approach tackles the point cloud registration by directly matching superpoints. The rich information in the matching step allows us to effectively filter out outliers without using RANSAC-like approaches. As a result, our approach is efficient and runs fast in practice.
  • Simplicity. Our registration pipeline can be trained end-to-end without the any ad hoc postprocessing refinement. It streamlines the process and makes it more straightforward to implement and understand.
  • Robustness. We thoroughly evaluate our method on various datasets, demonstrating its superior performance in comparison to existing techniques. Our approach achieves state-of-the-art results on multiple benchmarks, with notably high inlier ratios. This showcases the robustness and accuracy of our proposed registration framework.
2 Related Work
Traditional registration approaches. The most known algorithm Iterative Closest Point (ICP) [5] has been widely used for point cloud registration. ICP solves the registration problem iteratively in two steps: (1) It obtains the spatially closest point correspondence and then (2) finds the least-squares rigid transformation. The spatial-distance-based correspondences are sensitive to the initial transformation and point noises. A lot of variants [46, 11, 49, 6, 45, 35] have been proposed to improve ICP. Another line of work [27, 17] formulates the point cloud registration as probability distribution matching problems to improve the robustness and to converge quickly. However, they still heavily rely on an appropriate initialization and can easily converge to a local optimum as opposed to the global solution. Unlike the local methods mentioned above, global approaches are invariant to the initialization. Based on branch-and-bound techniques, several works [61, 8, 9, 35] have been proposed, but they are very slow and impractical in some scenarios. Another method is to extract and match keypoints based on feature extraction methods such as FPFH [47] and SHOT [52]. Then RANSAC [18] can be used for registration. However, RANSAC is computationally very slow compared to ICP.
Learning-based registration approaches. Recently, many works have used deep learning for point cloud learning and registration. Some work first estimates the correspondence between two point clouds and then computes the transformation with some robust pose estimators. To predict the correspondence between two point clouds, 3DMatch [68] detects the repeatable keypoints and learns discriminative descriptors for keypoints. The following works aim to either improve the keypoint detections [4, 32, 62] or learn better feature descriptors [13, 14, 15, 29, 1, 55]. Predator [24] uses the attention mechanism proposed in Transformers [54] to enhance the point feature descriptors. Other detector-free methods [66, 44] extract the correspondences by considering all possible matches. Another line of work [10, 67] has included the transformation computation into the training pipeline. Following the idea of ICP [5],  [57, 56, 63, 19, 31] iteratively predict the soft correspondences and computes the transformation with differentiable weighted SVD. Another way [2, 33, 25, 59, 48] first extracts a global feature vector for each point cloud and predicts the transformation with global feature vectors. This approach usually fails in large-scale scenes. Unlike these works which require eighter ad-hoc postprocessing or coarse-to-fine registration, our method directly matches the superpoints without any refinement.
Correspondence filters. RANSAC [18] is typically used to filter out the outliers in the predicted correspondence to obtain a robust transformation estimation. However, RANSAC is relatively slow and cannot be incorporated into the training pipeline because the hypothesis selection step is non-differentiable. To alleviate these problems, DSAC [7] modify the RANSAC pipeline and make it differentiable. Other deep robust estimators [3, 12, 40, 20, 30, 65] usually use the classification network to identify which correspondences are outliers and then reject them. Instead of using these complex correspondence filters, our method can directly filter out outliers effectively by leveraging the rich information in the superpoints matching.
3 Method
Given the source and target point clouds $\mathbf{X}\in\mathbb{R}^{M\times 3}$ and $\mathbf{Y}\in\mathbb{R}^{N\times 3}$, our goal is to determine the optimal rigid transformation $\mathbf{T}=\{\mathbf{R},\mathbf{t}\}$ with rotation $\mathbf{R}\in SO(3)$ and translation $\mathbf{t}\in\mathbb{R}^{3}$ to align two point clouds into a common coordinate system. $M$ and $N$ denote the numbers points.
3.1 Superpoints Feature Extraction and Enhancement
In our approach, we use Kernel Point Convolution (KPConv) [51] as the backbone to selectively downsample the point cloud into a set of superpoints and extract global feature vectors for each superpoint. Specifically, the KPConv backbone uses a series of ResNet-like blocks [23] and convolutions to downsample the input point clouds into a reduced set of superpoints $\mathbf{X}{{}^{\prime}}\in\mathbb{R}^{M^{\prime}\times 3}$ and $\mathbf{Y}{{}^{\prime}}\in\mathbb{R}^{N^{\prime}\times 3}$, where $M^{\prime}<M$ and $N^{\prime}<N$. The superpoints are described by their feature vectors $\mathbf{F}_{\mathbf{X}{{}^{\prime}}}\in\mathbb{R}^{M^{\prime}\times D}$ and $\mathbf{F}_{\mathbf{Y}{{}^{\prime}}}\in\mathbb{R}^{N^{\prime}\times D}$, respectively, with $D$ being the feature dimension. The network weights are shared among the two point clouds. We use a shallower backbone for 3DMatch dataset compared to [64, 24] to avoid significant downsampling.
Figure 2: Model Architecture: The KPConv backbone downsamples the input point cloud and generates superpoints and superpoint feature vectors. These superpoint features are then conditioned on the other point cloud in the Feature Enhancement block. Lastly, superpoint features can be directly matched using Global Softmax to predict correspondences and estimate rigid transformation.
Although KPConv backbone provides reasonably good representations, the superpoints features are obtained within each point cloud independently. To obtain highly discriminative feature representations for superpoints matching, we further enhance their feature representations in the source point cloud to be conditioned on the target one and vice-versa. Following the previous work of point cloud registration [64] and the 2D counterpart of superpoints matching (e.g., optical flow) [60], we adopt the multi-head attention mechanism in the Transformer model [54] as the feature enhancement module, shown in Fig. 2. It consists of both self and cross-attention, where the self-attention is to integrate the information from the other points within the same point cloud and the cross-attention allows interactions with points in another point cloud to consider the mutual dependencies. In addition to the multi-head attention, other components in the Transformer model, including position encodings of 3D points, residual connections, layer normalization, and feed-forward network are applied to each layer. The entire feature enhancement module consists of 6 such layers with 256 dimensions and 8 attention heads.
The outputs of the feature enhancement module are features $\bar{\mathbf{F}}_{\mathbf{X}{{}^{\prime}}}\in\mathbb{R}^{M^{\prime}\times D}$ and $\bar{\mathbf{F}}_{\mathbf{Y}{{}^{\prime}}}\in\mathbb{R}^{N^{\prime}\times D}$ which has aggregated geometric information from both source and target point cloud. The strongly associated features are strengthened while the weakly associated features are weakened. With such highly discriminative feature representations, we can obtain high-quality superpoints matchings.
3.2 Direct Superpoint Matching for Rigid Transformation Estimation
To get the correspondences between two point clouds, we first compare the feature similarity for each point in $\bar{\mathbf{F}}_{\mathbf{X}{{}^{\prime}}}$ to all points in $\bar{\mathbf{F}}_{\mathbf{Y}{{}^{\prime}}}$ by computing their correlations [60], which can be done efficiently in a single step as follows:
${\mathbf{C}}=\mathrm{softmax}\left(\bar{\mathbf{F}}_{\mathbf{X}{{}^{\prime}}}\bar{\mathbf{F}}_{\mathbf{Y}{{}^{\prime}}}^{T}\right)\in\mathbb{R}^{M^{\prime}\times N^{\prime}},$ (1)
where $\mathbf{C}$ is the normalized correlation matrix that represents the similarity between two point clouds. Based on the correlation matrix, the correspondences $\mathbf{\hat{Y}}$ and $\mathbf{\hat{X}}$ can be directly calculated by using the largest correlation for each point.
The rigid transformation between the source and target point clouds can be then estimated using the superpoints correspondences with a weighted variant of the Kabsch-Umeyama algorithm [28, 53]:
$\mathbf{\hat{R},\hat{t}}=\operatorname*{arg\,min}_{\mathbf{R},\mathbf{t}}\sum_{i}^{\min(N^{\prime},M^{\prime})}{w_{i}\lVert\mathbf{R}\mathbf{\hat{x}}_{i}+\mathbf{t}-\mathbf{\hat{y}}_{i}\rVert}^{2},$ (2)
where $\mathbf{\hat{x}}_{i}$, $\mathbf{\hat{y}}_{i}$ are two matched superpoints, denoting the $i^{th}$ point in $\mathbf{\hat{X}}$, $\mathbf{\hat{Y}}$, respectively. The coefficient $w_{i}$ can be used to weigh different correspondences.
Although we use the highly discriminative feature representations enhanced by the attention module, the correspondences are inevitably noisy. How to filter out the outliers (i.e., incorrect correspondences)? We show that the normalized correlation matrix $\mathbf{C}$ obtained from superpoints matching in Eq. (1) contains rich information, which allows us effectively select high-confident superpoints correspondences to estimate the transformation. Specifically, if $\hat{\mathbf{x}}_{i}$ is similar to multiple superpoints, e.g., because of the repetitive patterns, its matching to $\hat{\mathbf{y}}_{i}$ tends to be unreliable. Therefore, the normalized correlation score between them will be low since $\mathbf{C}$ is normalized w.r.t. all other superpoints in the target point cloud. We can therefore set $w_{i}=\mathbf{C}(\hat{\mathbf{x}}_{i},\hat{\mathbf{y}}_{i})$, capturing the confidence of the superpoints matchings. On the other hand, we can use $w_{i}$ to select only highly confident superpoint matchings, e.g., top 15% of them, to estimate the transformation in Eq.(2. We can also augment the weights $w_{i}$ by considering the overlap score, which is introduced below, to further improve the accuracy.
Compared to previous work [44, 66, 38], our approach is simple yet effective, which does not only eliminates the coarse-to-fine strategy but more importantly, the slow and non-differentiable RANSAC. As a result, our approach is fast and the superpoints feature learning, superpoints matching, and transformation estimation can be optimized jointly in an end-to-end manner, leading to superior accuracy. Although the feature enhancement module is also used in [64], our approach is fundamentally different. Instead of predicting the correspondences by using the feature representations of the source point cloud only, we perform superpoints matching and use the rich information in the matching step to effectively filter outliers and get a robust estimation of the transformation.
3.3 Loss Functions
We train our approach using the following three loss functions, where the transformation loss is the main loss term and the other two are auxiliary ones.
Transformation Loss. We apply a L1 loss on the predicted transformed locations of all keypoints with the predicted and ground truth transformation matrix.
$\mathcal{L}_{T}=\frac{1}{M^{\prime}}{\sum_{i}^{M^{\prime}}}\left\lvert\hat{\mathbf{R}}\mathbf{x_{i}^{\prime}}+\hat{\mathbf{t}}-\left(\mathbf{R}_{gt}\mathbf{x_{i}^{\prime}}+\mathbf{t}_{gt}\right)\right\rvert_{1}.$ (3)
Overlap Loss. Inspired by [64], we estimate the overlap values $\mathbf{\hat{O}_{\mathbf{X^{\prime}}}}$ and $\mathbf{\hat{O}_{\mathbf{Y^{\prime}}}}$ using a separate MLP layer based on the enhanced feature $\bar{\mathbf{F}}_{\mathbf{X}{{}^{\prime}}}$ and $\bar{\mathbf{F}}_{\mathbf{Y}{{}^{\prime}}}$, respectively. The overlap estimation is formulated as a binary classification problem, so we use binary cross-entropy loss:
$\mathcal{L}^{X}_{o}=-\frac{1}{M^{\prime}}\sum_{i}^{M^{\prime}}{o_{x,i}^{*}\cdot\log\hat{o}_{x,i}^{*}+(1-o_{x,i}^{*})\cdot\log\left(1-\hat{o}_{x,i}\right)},$ (4)
where $\hat{o}_{x,i}$ is the estimated overlap probability and $o_{x,i}^{*}$ is the ground truth label. We compute the overlap loss $\mathcal{L}^{Y}_{o}$ for the target point cloud similarly.
Feature Loss. Following [64], to ensure that the enhanced features of both point clouds are in the same feature space, we apply an InfoNCE [39] loss on the enhanced features $\bar{\mathbf{F}}_{\mathbf{X}{{}^{\prime}}}$ and $\bar{\mathbf{F}}_{\mathbf{Y}{{}^{\prime}}}$. Given a set of superpoints correspondences $\{(\hat{\mathbf{x}}_{i},\hat{\mathbf{y}}_{i})\}_{i=1}^{K}$ and their associated feature representations $\{(\bar{\mathbf{f}}_{\hat{\mathbf{x}}_{i}},\bar{\mathbf{f}}_{\hat{\mathbf{y}}_{i}})$, the feature loss is defined as
$\mathcal{L}_{f}=-\frac{1}{K}\log\frac{\bar{\mathbf{f}}_{\hat{\mathbf{x}}_{i}}^{T}\mathbf{W}\bar{\mathbf{f}}_{\hat{\mathbf{y}}_{i}}}{\bar{\mathbf{f}}_{\hat{\mathbf{x}}_{i}}^{T}\mathbf{W}\bar{\mathbf{f}}_{\hat{\mathbf{y}}_{i}}+\sum_{j\neq i}\bar{\mathbf{f}}_{\hat{\mathbf{x}}_{i}}^{T}\mathbf{W}\bar{\mathbf{f}}_{\hat{\mathbf{y}}_{j}}}.$ (5)
The linear transformation $\mathbf{W}$ is enforced to be symmetrical by parameterizing it as the sum of an upper triangular matrix $\mathbf{U}$ and its transpose, i.e. $\mathbf{W}=\mathbf{U}+\mathbf{U}^{T}$.
The final loss is a weighted sum of all the losses above with
$\mathcal{L}=\mathcal{L}_{T}+\alpha\mathcal{L}_{f}+\beta(\mathcal{L}_{o}^{X}+\mathcal{L}_{o}^{Y}),$ (6)
where we set the loss weights $\alpha=0.1$ and $\beta=1$ empirically.
4 Experiments
We evaluate our approach on two benchmarks with overlap ranging from 10% to 75%. The first dataset we evaluate is on ModelNet with two benchmarks settings following [24, 64] and the second dataset is 3DMatch [68] on two benchmarks following [24, 64, 44, 66].
4.1 Implementation details
Our approach is implemented using the PyTorch framework [41] on a system with an Intel i9-1300K CPU and a single RTX 3090 GPU. The network training is performed with the AdamW optimizer [36], utilizing a learning rate of 0.0001 and a weight decay of 0.0001. For the ModelNet dataset, the network is trained for 400 epochs with a batch size of 4. On the 3DMatch dataset, the network is trained for 50 epochs with a batch size of 4 as well. Training the network requires approximately 22 hours on ModelNet and around 2 days on 3DMatch.
4.2 ModelNet and ModelLoNet Benchmarks
The ModelNet40 [58] dataset comprises of synthetic CAD models. Following the data setting in [24, 64], the point clouds are randomly sampled from mesh faces of the CAD models, cropped and subsampled. For the ModelNet and ModelLoNet benchmarks, the average overlap is 73.5% and 53.6%, respectively. Our network is trained exclusively on ModelNet, and we evaluate its generalization performance on ModelLoNet. For benchmarking the performance of our model we use the Relative Rotation Error (RRE) and Relative translation Error (RTE) as the primary metrics. Following [64], we also calculate the Chamfer Distance (CD) between the registered scan pairs.
Table 1: Registration results on the ModelNet and ModelLoNet benchmarks.
Methods ModelNet ModelLoNet
[…]
PointNetLK 29.725 0.297 0.0235 48.567 0.507 0.0367
OMNet 2.947 0.032 0.0015 6.517 0.129 0.0074
DCP-v2 11.975 0.171 0.0117 16.501 0.300 0.0268
RPM-Net 1.712 0.018 0.00085 7.342 0.124 0.0050
Predator 1.739 0.019 0.00089 5.235 0.132 0.0083
RegTR 1.473 0.014 0.00078 3.930 0.087 0.0037
Geotransformer 1.568 0.018 - 3.809 0.102 -
Ours 1.247 0.011 0.00074 3.809 0.088 0.0040
The results are shown in Table 1. We compare against correspondence-based approaches [24, 64], coarse-to-fine registration approaches [44], and end-to-end methods [2, 63, 59]. We see that our approach performs well on both benchmarks. Moreover, the low chamfer error suggests that predicted correspondences have very high accuracy. Our approach is also able to outperform methods using post-processing steps like RANSAC [24] by a significant margin, which further strengthens the point that direct superpoints matchings can actually work well for point cloud registration.
4.3 3DMatch and 3DLoMatch Benchmarks
3DMatch [68] is a collection of 62 scenes, from which we use 46 for training, 8 for testing, and 8 for validation following [64, 24, 44]. We use the preprocessed data from [24] which contains point clouds downsampled using a voxel-grid subsampling method. The 3DMatch benchmark contains point clouds pairs with >30% overlap while the 3DLoMatch benchmark contains scan pairs with only 10%-30% overlap. Following [64], we perform training data augmentation by applying small rigid perturbations, jittering, and shuffling of points. Following the literature [24, 44, 68], We report the results of 3DMatch dataset on 5 metrics including RRE, RTE, Registration Recall (RR), Feature Matching Recall (FMR), and Inlier Ratio (IR).
We compare our approach against several approaches including learned corresponding based algorithms [21, 13, 4, 24] and coarse-to-fine approaches [66, 44]. Since the algorithms which use some post-preprocessing steps like RANSAC perform well only with a large number of interest points, we only show results for the maximum number of sampled points (5000). We also compare against several approaches designed to avoid RANSAC [10, 59, 12].
Table 2: Registration results on the 3DMatch and 3DLoMatch benchmarks.
Methods 3DMatch 3DLoMatch
[…]
3DSN 2.199 0.071 78.4 - - 3.528 0.103 33.0 - -
FCGF 1.949 0.066 85.1 97.4 56.8 3.147 0.100 40.1 76.6 21.4
D3Feat 2.161 0.067 81.6 95.6 39.0 3.361 0.103 37.2 67.3 13.2
Predator 2.029 0.064 89.0 96.6 58.8 3.048 0.093 59.8 78.6 26.7
OMNet 4.166 0.105 35.9 - - 7.299 0.151 8.4 - -
DGR 2.103 0.067 85.3 - - 3.954 0.113 48.7 - -
PCAM 1.808 0.059 85.5 - - 3.529 0.099 54.9 - -
RegTR 1.567 0.049 92.0 - - 2.827 0.077 64.8 - -
YOHO - - 90.8 98.2 64.4 - - 65.2 79.4 25.9
CofiNet - - 89.3 98.1 49.8 - - 67.5 83.1 24.4
Geo-Transformer (SVD) - - 86.5 - - - - 59.9 - -
GeoTransformer (LGR) 1.625 0.053 91.5 97.7 70.3 2.547 0.074 74.0 88.1 43.3
Ours 1.436 0.045 93.7 96.5 89.8 2.553 0.074 65.0 76.5 57.5
The quantitative results are shown in Table 2. For the 3DMatch benchmark, our approach outperforms all the previous methods. This implies that in cases of significant overlap (>30%), the superpoint correspondences are very distinctive and accurate. One of the problems associated with superpoint matching in point clouds is the resolution issue, where the correspondences might not actually coincide with each other when transformed with the ground truth transformation matrix due to the subsampling of the point cloud. In our case, the resolution issue is automatically handled by the correlation weights. We also show systematic comparisons with other approaches in Table 3. We can see our model runs fast and is compact.
Table 3: Comparisons of inference speed and number of parameters on the 3DMatch dataset.
Methods Matching e2e Time Param
CoFiNet coarse-to-fine 1.922s 5.5M
GeoTransformer coarse-to-fine 0.088s 9.8M
RegTR 0.063s 11.8M
Ours direct 0.073s 7.8M
For the 3DLoMatch benchmark, in comparison to all the approaches that do not use any post-processing, our method performs significantly better than the other approaches. Note that [44] uses LGR as the refinement step to get the best results, but when using all the predicted correspondences with their scores and applying weighted SVD, they get much worse results. This implies that the final set of all the correspondences is not optimal and contains outliers. In our case, these outliers are filtered by the overlap and correlation values, thus providing accurate correspondences. The validity of the correspondences predicted by our approach can be verified by comparing the mean Inlier Ratio in Table 2. We get the highest mean IR, almost 19% better than the second-best approach on 3DMatch and 14% higher in 3DLoMatch.
4.4 Ablation Studies
Table 4: Comparisons of different outlier filtering methods.
Outlier Filtering Method 3DMatch 3DLoMatch
RRE RTE RR RRE RTE RR Time
No filtering 1.462 0.046 93.3 2.652 0.074 64.6 0.073
Ours + RANSAC-50k 2.788 0.090 82.3 4.653 0.115 32.0 0.141
Ours (top 15%) + RANSAC-50k 1.701 0.051 93.0 2.696 0.078 64.5 0.153
Iterative refinement 1.404 0.043 93.8 2.592 0.071 65.7 0.078
Correlation scores (top 15%) 1.436 0.045 93.7 2.553 0.074 65.0 0.073
Refer to caption Refer to caption
Figure 3: Left: Effect of increasing the number of attention layers. Right: Comparisons of outlier filtering strategies.
Effectiveness of the feature enhancement. The feature enhancement module is essential in our approach. To assess its effectiveness, we trained various networks with attention layers ranging from 0 to 6. Figure 3 illustrates the performance changes as we increase the number of attention layers. When no feature enhancement is used, the model’s performance is significantly poorer, which is understandable since the source point cloud’s superpoint features lack information about the target point cloud. Performance stabilizes at around 6 attention layers. It is worth noting that increasing the number of layers also leads to higher computation burden.
Using correlation scores for filtering out outliers. To filter out outliers in the predicted correspondences we primarily use the correlation values along with overlap values as weights for weighted SVD. Since these weights are much lower for outlier values, the registration results are very accurate. If we use unweighted SVD (with equal weights for all the correspondences), the performance drops rapidly with the increase in number of correspondences, as shown in Fig. 1.
We experiment with a simple iterative refinement scheme proposed in [44], where we can iteratively re-estimate the transformation by pruning out outliers. We found that the performance of this refinement saturated after 5 iterations. See Table 4 row 4 for results. Fig. 1 also shows that this approach is quite stable in pruning outliers on different number of correspondences and improves the results by a slight margin.
Table 5: Effectiveness of different loss terms.
Overlap Feature 3DMatch 3DLoMatch
Loss Loss RRE RTE RR RRE RTE RR
2.521 0.076 76.5 5.272 0.132 31.2
2.123 0.062 79.6 4.020 0.105 37.2
1.651 0.049 90.3 2.876 0.082 58.0
1.436 0.045 93.7 2.553 0.074 65.0
We also experiment with RANSAC on the set of superpoints correspondences, See Table 4 row 2, 3. But as expected, RANSAC gives slightly worse results while taking more computation time. This implies that using our approach is much more effective in pruning outliers thus eliminating the need for refinement.
Effectiveness of the loss terms. Lastly, we analyze the effectiveness of each loss function. Table 5 shows the results with different loss function configurations. We see that by just using Feature Loss with Transformation loss, the model is able to achieves good performance and using Overlap loss is the final step that helps the model prune remaining outliers. Row 3 in Table 5 again implies that superpoint correspondences are very distinctive.
Figure 4: Qualitative Results on 3Dmatch dataset: The source cloud is shown in Green with the red color representing the matched superpoints. Similarily target cloud is in red color with matching superpoints in green color.
5 Limitations and Broader Impacts
We have identified two limitations in the current architecture, First, our approach requires a decent amount of superpoints in the low overlapping region to perform well. Second, our approach samples about 200 best correspondences, but for the low overlapping region only a few of them lie on the overlapping region thus making it hard for the network to achieve good performance on 3DLoMatch benchmark. Instead of regularly downsampling, structure-aware sampling strategies are needed.
This research has the potential to significantly impact the fields of 3D computer vision and robotics, as well as any domains that rely on accurate and efficient point cloud registration, such as autonomous navigation, 3D mapping, and augmented reality. By introducing a simple yet effective approach for directly matching superpoints in an end-to-end manner, the research not only speeds up the registration process but also enhances its robustness by effectively filtering out outliers without any postprocessing refinement.
6 Conclusion
We introduce a straightforward yet highly effective method for point cloud registration that offers simplicity in understanding and implementation. Our approach directly matches superpoints to establish correspondences, enabling the computation of rigid transformations and correlation weights. To enhance performance in low overlap regions, we augment the correlation weights with predicted overlap values. Experimental evaluations on the 3DMatch and ModelNet datasets demonstrate the efficacy of our approach across different overlap scenarios. Notably, our approach is computationally efficient, robust to outliers, and does not require post-processing steps, making it suitable for real-time applications. In the future, we plan to extend its application to cross-modalities such as image-to-image and image-to-point cloud registration.
Appendix A Appendix
A.1 Network Architecture
Figure 5: Network Architecture of the KPConv backbone for 3DMatch and ModelNet dataset. We do not downsample the point cloud by a large factor to maintain resolution on the overlapping region of the point cloud pair.
Figure 6: Network architecture of the Feature Enhancement Module. The architecture is shown for just one block of self-attention and cross-attention layers but can be extended similarly for improvement in performance.
The network architecture for the KPConv backbone used is shown in Fig. 5 and the architecture for the Feature Enhancement module is shown in Fig. 6.
We do not downsample the point cloud by a large factor in the bakbone unlike other methods [44, 64] as this leads to a very low resolution in the overlapping region of the point cloud pair. Since, we are training the network to learn correlations between the downsampled superpoints, it is beneficial to have a moderately larger number of points.
A.2 Evaluation Metrics
We report Registration Recall (RR), Relative Rotation Error (RRE), Relative Translation Error (RTE), Feature Matching Recall (FMR), and Inlier Ratio (IR) for the 3DMatch dataset following common practice [44, 64, 24, 66, 4]. For ModelNet dataset, we report RRE, RTE, and Chamfer Distance Error (CD) following [64, 24, 59].
A.2.1 3DMatch/3DLoMatch
Inlier Ratio is the fraction of inlier correspondences in all predicted correspondences. Corresponding points are considered inliers if the distance between them under the ground-truth transformation is smaller than $\tau_{1}=10cm$
[…]
Feature Matching Recall is the fraction of point cloud pairs whose inlier ratio is above $\tau_{2}=0.05$. It indicates the likelihood of recovering the optimal transformation between two point clouds. It is a good indicator for methods which use post-processing refinement steps like RANSAC [66, 44, 24] as having a higher FMR increases the probability of recovering the correct transformation. If the total number of point cloud pairs in the test set is $\mathcal{N}$, FMR can be calculated as
$\textnormal{FMR}=\frac{1}{\mathcal{N}}\sum^{\mathcal{N}}_{i=1}\llbracket\textnormal{IR}>\tau_{2}\rrbracket.$ (8)
Registration Recall represents the fraction of point cloud pairs whose Root Mean Square Error (RMSE) is less than $\tau_{3}=20cm$. RR directly measures the effectiveness of the registration algorithm.
$\textnormal{RR}=\frac{1}{\mathcal{N}}\sum^{\mathcal{N}}_{i=1}\left\llbracket\sqrt{\frac{1}{|\mathcal{C^{*}}|}\sum_{(\mathbf{x}_{i},\mathbf{y}_{i})\in\mathcal{C^{*}}}\lVert\mathbf{T}^{\mathbf{X}}_{\mathbf{Y}}(\mathbf{x}_{i})-\mathbf{y}_{i}\rVert^{2}_{2}}<\tau_{3}\right\rrbracket,$ (9)
where $\mathcal{C^{*}}$ is the set of ground truth correspondences and $\mathbf{T}^{\mathbf{X}}_{\mathbf{Y}}$ is the predicted transformation matrix.
Relative Rotation Error and Relative Translation Error are the rotation and translation errors between the ground truth transformation matrix $\mkern 1.5mu\overline{\mkern-1.5mu\mathbf{T}\mkern-1.5mu}\mkern 1.5mu^{\mathbf{X}}_{\mathbf{Y}}\in SE(3)$ and the predicted transformation $\mathbf{T}^{\mathbf{X}}_{\mathbf{Y}}\in SE(3)$. We can obtain $\mathbf{R}\in SO(3)$, $\mkern 1.5mu\overline{\mkern-1.5mu\mathbf{R}\mkern-1.5mu}\mkern 1.5mu\in SO(3)$, $\mathbf{t}\in\mathbf{R}^{3}$ and $\mkern 1.5mu\overline{\mkern-1.5mu\mathbf{t}\mkern-1.5mu}\mkern 1.5mu\in\mathbf{R}^{3}$ as the predicted rotation matrix, ground truth rotation matrix, predicted translation vector and ground truth translation vector respectively. Thus, RRE and RTE can be calculated as follows
$\textnormal{RRE}=\textnormal{arccos}\left(\frac{\textnormal{trace}(\mathbf{R}^{T}\mkern 1.5mu\overline{\mkern-1.5mu\mathbf{R}\mkern-1.5mu}\mkern 1.5mu)-1}{2}\right),~{}~{}~{}\textnormal{RTE}=\lVert\mathbf{t}-\mkern 1.5mu\overline{\mkern-1.5mu\mathbf{t}\mkern-1.5mu}\mkern 1.5mu\rVert.$ (10)
A.2.2 ModelNet/ModelLoNet
Chamfer Distance measures the dissimilarity or discrepancy between two sets of points or point clouds. It quantifies the minimum average distance between each point in one set and its nearest neighbor in the other set. It provides a measure of how well two point clouds align or match with each other, with lower values indicating a better alignment.
$\text{{CD}}(\mathbf{X},\mathbf{Y})=\sum_{x\in\mathbf{X}}\min_{y\in\mathbf{Y}}\|y-\mathbf{T}^{\mathbf{X}}_{\mathbf{Y}}(x)\|_{2}^{2}+\sum_{y\in\mathbf{Y}}\min_{x\in\mathbf{X}}\|y-\mathbf{T}^{\mathbf{X}}_{\mathbf{Y}}(x)\|_{2}^{2},$ (11)
where $\mathbf{X}$ and $\mathbf{Y}$ represent the sets of points in two point clouds, $\|\cdot\|_{2}$ denotes the Euclidean distance and $\mathbf{T}^{\mathbf{X}}_{\mathbf{Y}}$ is the predicted transformation between point clouds $\mathbf{X}$, and $\mathbf{Y}$.
A.3 Additional Results
Table 6 shows the scene-wise registration results on the 3DMatch and 3DLoMatch benchmarks. Following literature [44, 64, 24], we only calculate mean RRE and RTE for the successfully registered scan pairs. For the 3DMatch benchmark, our method outperforms other methods in every metric. Our methods achieves not only the highest registration recall, but also the lowest mean RRE and RTE. This implies that the for all the successfully registered points clouds, the predicted correspondences and their correlation values are very accurate. For the 3DLoMatch benchmark, our method outperforms every method which does not use post-processing refinement like RegTR, while giving the lowest RRE and RTE values. In comparison to approaches which use post-processing refinement like GeoTransformer, Predator, and CoFiNet, the performance gap is expected as the predictions in 3DLoMatch contain a much higher number of outliers. But the gap is minor and our approach runs faster.
Table 6: Scene-wise registration results on the 3DMatch and 3DLoMatch benchmarks.
Method 3DMatch 3DLoMatch
Kitchen Home 1 Home 2 Hotel 1 Hotel 2 Hotel 3 Study Lab Mean Kitchen Home 1 Home 2 Hotel 1 Hotel 2 Hotel 3 Study Lab Mean
Registration Recall (%)normal-$\uparrow$
3DSN 90.6 90.6 65.4 89.6 82.1 80.8 68.4 60.0 78.4 51.4 25.9 44.1 41.1 30.7 36.6 14.0 20.3 33.0
FCGF 98.0 94.3 68.6 96.7 91.0 84.6 76.1 71.1 85.1 60.8 42.2 53.6 53.1 38.0 26.8 16.1 30.4 40.1
D3Feat 96.0 86.8 67.3 90.7 88.5 80.8 78.2 64.4 81.6 49.7 37.2 47.3 47.8 36.5 31.7 15.7 31.9 37.2
Predator 97.6 97.2 74.8 98.9 96.2 88.5 85.9 73.3 89.0 71.5 58.2 60.8 77.5 64.2 61.0 45.8 39.1 59.8
CoFiNet 96.4 99.1 73.6 95.6 91.0 84.6 89.7 84.4 89.3 76.7 66.7 64.0 81.3 65.0 63.4 53.4 69.6 67.5
RegTR 97.8 90.6 75.5 97.8 94.9 100 88.5 91.1 92.0 66.2 58.5 64.9 72.7 61.3 70.7 53.0 71.0 64.8
GeoT 98.9 97.2 81.1 98.9 89.7 88.5 88.9 88.9 91.5 85.9 73.5 72.5 89.5 73.2 66.7 55.3 75.7 74.0
Ours 97.6 88.7 87.4 97.8 93.6 100 87.2 97.8 93.7 62.3 55.3 69.4 72.2 62.0 45.6 46.2 76.8 65.0
Relative Rotation Error (°)normal-$\downarrow$
3DSN 1.926 1.843 2.324 2.041 1.952 2.908 2.296 2.301 2.199 3.020 3.898 3.427 3.196 3.217 3.328 4.325 3.814 3.528
FCGF 1.767 1.849 2.210 1.867 1.667 2.417 2.024 1.792 1.949 2.904 3.229 3.277 2.768 2.801 2.822 3.372 4.006 3.147
D3Feat 2.016 2.029 2.425 1.990 1.967 2.400 2.346 2.115 2.161 3.226 3.492 3.373 3.330 3.165 2.972 3.708 3.619 3.361
Predator 1.861 1.806 2.473 2.045 1.600 2.458 2.067 1.926 2.029 3.079 2.637 3.220 2.694 2.907 3.390 3.046 3.412 3.048
CoFiNet 1.910 1.835 2.316 1.767 1.753 1.639 2.527 2.345 2.011 3.213 3.119 3.711 2.842 2.897 3.194 4.126 3.138 3.280
RegTR 1.729 1.34 1.79 1.639 1.28 1.810 1.57 1.35 1.56 3.366 2.446 3.244 2.732 2.439 2.919 3.044 2.428 2.827
GeoT 1.79 1.353 1.79 1.52 1.328 1.57 1.952 1.678 1.625 2.35 2.30 2.541 2.455 2.49 2.504 3.01 2.716 2.547
Ours 1.780 1.272 1.596 1.457 1.239 1.456 1.492 1.196 1.436 2.330 2.281 2.59 2.72 2.520 2.58 2.775 2.61 2.55
Relative Translation Error (m)normal-$\downarrow$
3DSN 0.059 0.070 0.079 0.065 0.074 0.062 0.093 0.065 0.071 0.082 0.098 0.096 0.101 0.080 0.089 0.158 0.120 0.103
FCGF 0.053 0.056 0.071 0.062 0.061 0.055 0.082 0.090 0.066 0.084 0.097 0.076 0.101 0.084 0.077 0.144 0.140 0.100
D3Feat 0.055 0.065 0.080 0.064 0.078 0.049 0.083 0.064 0.067 0.088 0.101 0.086 0.099 0.092 0.075 0.146 0.135 0.103
Predator 0.048 0.055 0.070 0.073 0.060 0.065 0.080 0.063 0.064 0.081 0.080 0.084 0.099 0.096 0.077 0.101 0.130 0.093
CoFiNet 0.047 0.059 0.063 0.063 0.058 0.044 0.087 0.075 0.062 0.080 0.078 0.078 0.099 0.086 0.077 0.131 0.123 0.094
RegTR 0.040 0.041 0.057 0.057 0.042 0.039 0.054 0.058 0.049 0.079 0.064 0.078 0.094 0.074 0.060 0.093 0.077 0.077
GeoT 0.042 0.046 0.059 0.055 0.046 0.050 0.073 0.053 0.053 0.062 0.070 0.071 0.080 0.075 0.049 0.107 0.083 0.074
Ours 0.043 0.037 0.061 0.051 0.044 0.029 0.048 0.044 0.045 0.061 0.068 0.068 0.088 0.071 0.054 0.091 0.091 0.074
A.4 Iterative Pose Refinement
We experiment with a simple iterative refinement scheme proposed in [44], where we can iteratively re-estimate the transformation by pruning out outliers.
$\mathbf{R,t}=\max_{\mathbf{R_{i}},\mathbf{t_{i}}}\sum_{i}{\llbracket\lVert\mathbf{R}\mathbf{x}_{i}+\mathbf{t}-\mathbf{y}_{i}\rVert^{2}<\tau_{a}\rrbracket},$ (12)
[…]
Refer to caption Refer to caption
Figure 7: Improvement in Reigstration Recall on increasing the number of iterations for pose refinement. We see that the performance saturates after the 5th iteration for 3DLoMatch and after 2 iterations for 3DMatch.
A.5 Qualitative Results
We provide more visual results in Fig. 8 and Fig. 9.
Figure 8: Qualitative Results on the ModelLoNet dataset. Row (a)-(d) show the cases with successful registration. While the rows (e) and (f) show the failure cases. The source cloud is shown in Green with the red color representing for the matched superpoints. Similarily target cloud is in red color with matching superpoints in green color. We see that our approach only fails in cases which are very ambigous.
Figure 9: Qualitative Results on the 3DLoMatch dataset. Row (a)-(d) show the cases with successful registration. While the rows (e) and (f) show the failure cases. The source cloud is shown in Green with the red color representing for the matched superpoints. Similarily target cloud is in red color with matching superpoints in green color.

### v4

@@
-Direct Superpoints Matching for Fast and Robust Point Cloud Registration
+Direct Superpoints Matching for Robust Point Cloud Registration
 Abstract
-Although deep neural networks endow the downsampled superpoints with discriminative feature representations, directly matching them is usually not used alone in state-of-the-art methods, mainly for two reasons.
-First, the correspondences are inevitably noisy, so RANSAC-like refinement is usually adopted.
-Such ad hoc postprocessing, however, is slow and not differentiable, which can not be jointly optimized with feature learning.
-Second, superpoints are sparse and thus more RANSAC iterations are needed.
-Existing approaches use the coarse-to-fine strategy to propagate the superpoints correspondences to the point level, which are not discriminative enough and further necessitates the postprocessing refinement.
-In this paper, we present a simple yet effective approach to extract correspondences by directly matching superpoints using a global softmax layer in an end-to-end manner, which are used to determine the rigid transformation between the source and target point cloud.
-Compared with methods that directly predict corresponding points, by leveraging the rich information from the superpoints matchings, we can obtain more accurate estimation of the transformation and effectively filter out outliers without any postprocessing refinement.
-As a result, our approach is not only fast, but also achieves state-of-the-art results on the challenging ModelNet and 3DMatch benchmarks.
-Our code and model weights will be publicly released.
-1 Introduction
+Deep neural networks endow the downsampled superpoints with highly discriminative feature representations.
+Previous dominant point cloud registration approaches match these feature representations as the first step, e.g., using the Sinkhorn algorithm.
+A RANSAC-like method is then usually adopted as a post-processing refinement to filter the outliers.
+Other dominant method is to directly predict the superpoint matchings using learned MLP layers.
+Both of them have drawbacks: RANSAC-based methods are computationally intensive and prediction-based methods suffer from outputing non-existing points in the point cloud.
+In this paper, we propose a straightforward and effective baseline to find correspondences of superpoints in a global matching manner.
+We employ the normalized matching scores as weights for each correspondence, allowing us to reject the outliers and further weigh the rest inliers when fitting the transformation matrix without relying on the cumbersome RANSAC.
+Moreover, the entire model can be trained in an end-to-end fashion, leading to better accuracy.
+Our simple yet effective baseline shows comparable or even better results than state-of-the-art methods on three datasets including ModelNet, 3DMatch, and KITTI.
+We do not advocate our approach to be the solution for point cloud registration but use the results to emphasize the role of matching strategy for point cloud registration.
+The code and models are available at https://github.com/neu-vi/Superpoints_Registration
+I Introduction
 Point cloud registration refers to the task of aligning two partially overlapping point clouds into a shared coordinate system.
-In this paper, we tackle the problem of rigid registration where the goal is to determine the optimal transformation matrix, including rotation and translation, from one point cloud (source) to the other (target).
+In this paper, we tackle this problem where the goal is to determine the transformation matrix, including rotation and translation, from one point cloud (source) to the other (target).
 It has attracted a lot of research interest due to its broad applications in SLAM (Simultaneous Localization and Mapping) [14, 44], autonomous driving [28, 32], 3D reconstruction [22, 19], etc.
-Figure 1: Our approach focuses on directly matching the predicted superpoints by computing correlation between them.
-The bold dots show the superpoints with highest correlation which are used as correspondences to correctly register the point cloud pair.
-A prevailing paradigm to solve the registration task is to leverage the correspondences of superpoints across two point clouds, which can be obtained either using separate keypoint detectors [1, 4, 13, 24] to capture salient and distinctive points or regions within a point cloud.
-With the rapid advancement in representation learning in point clouds using deep neural networks [42, 43, 51], the keypoint-free approaches have gained significant attention, where the point cloud is downsampled into superpoints.
-Although they may not be designed to capture the salient keypoints in the point cloud, the learned representations endow the superpoints with discriminative power so they can be matched across the source and target point clouds.
-The transformation matrix can then be obtained based on the superpoints correspondences using the Kabsh-Umeyama Algorithm [53, 28].
-Direct superpoint matching is typically not used by itself in existing approaches for point cloud registration for two reasons.
-First, the correspondences of superpoints inevitably contain errors, either because the point clouds contain noisy sensory data or because some of the correspondences are simply incorrect.
-As a result, postprocessing refinement is usually needed, for instance, using RANSAC to prune out the outliers [13, 14, 68].
-However, RANSAC is inherently slow and not differentiable, and thus cannot be integrated into the training step.
-Second, superpoints are essentially sparse, particularly because of downsampling in neural networks.
-Therefore, more RANSAC iterations are needed as the number of inliers (i.e., correct correspondences) is small, which imposes a significant computation burden.
-To this end, the coarse-to-fine registration scheme is usually adopted in state-of-the-art approaches [44, 66, 38], where superpoints correspondences serve as coarse correspondences only.
-They are then propagated to the point level, forming finer-level correspondences.
-Since point-wise features only capture local information, they lack enough discriminative power.
-Consequently, the noise in the point-wise correspondences is more prominent, which further necessitates the additional refinement step.
-Although the Local to Global Registration (LGR) module introduced in [44] improves the speed, which works similarly to RANSAC in an iterative manner, the selection and refinement of the transformation still remain non-differentiable.
-The entire model thus can not be optimized jointly, leading to inferior accuracy.
-In this paper, we present a simple yet effective approach for rigid point cloud registration by directly matching the superpoints.
-Specifically, superpoints features are first obtained from the KPConv backbone [51], which are then enhanced using Transformer blocks with interleaved multi-head self and cross attention modules [64], enabling effective learning of the structural information of from the target to the source point cloud.
-The enhanced superpoint features are finally matched using a Global Softmax layer inspired by the 2D matching work [60], which generates a correlation matrix capturing the pairwise relationships between the keypoints from the source and target point clouds.
-But unlike [60], we fully harness the rich information in the correlation matrix.
-We employ a differentiable variant of the Kabsh-Umeyama Algorithm [53, 28], considering the strengths of matchings, where the weights of the SVD (Singular Value Decomposition) are the correlation scores between two corresponding superpoints.
-By examining the correlation scores of the superpoints, we can easily identify the highly reliable correspondences without resorting to the time-consuming RANSAC or the non-differentiable LGR module [44].
-Our approach is fully differentiable and can be trained in an end-to-end manner, which enables joint optimization of the feature representation learning, superpoints matching, and transformation estimation, leading to improved alignment results.
-Moreover, our approach does not require ad-hoc postprocessing, e.g., RANSAC, it runs fast, making it potentially more useful in practice.
-Compared with the recent end-to-end method [64], our approach directly matches the superpoints.
-With the rich information obtained from the matching step, our approach can get a more accurate estimation of the transformation between the point clouds robustly by filtering out the outliers effectively.
-We run experiments on standard benchmarks, including ModelNet [58] and 3DMatch [68], and achieve state-of-the-art results.
-Extensive ablation studies validate the effectiveness of each module in our proposed approach.
-The key contributions of the paper can be summarized as follows:
-  • Efficiency.
-Our approach tackles the point cloud registration by directly matching superpoints.
-The rich information in the matching step allows us to effectively filter out outliers without using RANSAC-like approaches.
-As a result, our approach is efficient and runs fast in practice.
-  • Simplicity.
-Our registration pipeline can be trained end-to-end without the any ad hoc postprocessing refinement.
-It streamlines the process and makes it more straightforward to implement and understand.
-  • Robustness.
-We thoroughly evaluate our method on various datasets, demonstrating its superior performance in comparison to existing techniques.
-Our approach achieves state-of-the-art results on multiple benchmarks, with notably high inlier ratios.
-This showcases the robustness and accuracy of our proposed registration framework.
-2 Related Work
+Figure 1: Our approach directly matches the superpoints between the input point clouds to register them by estimating the $SE(3)$ transformation matrix.
+The correlation weights obtained from matching are used to filter out incorrect correspondences (outliers) and further weigh the rest inliers for the transformation matrix estimation.
+The source cloud is shown in blue, with superpoints highlighted in red.
+Similarly, the target point cloud is displayed in yellow, with matching superpoints shown in blue.
+The intensity of the red and blue colors represents the correspondence weights.
+Notice how most keypoints are distributed around the three-way junction and exhibit high correspondence weights.
+(Best viewed in color.)
+A prevailing paradigm to solve the registration task is to leverage the correspondences of superpoints across two point clouds, which can be obtained using either keypoint detectors [1, 4, 11, 21] or downsampling in deep neural networks [37, 38, 47].
+Ideally, these superpoints should capture salient and distinctive points or regions within a point cloud.
+With the learned feature representations, the superpoints are endowed with sufficient discriminative power so they can be matched across the source and target point clouds, where the matching strategy is crucial for finding correspondences of superpoints and can significantly affect the convergence and performance of the network.
+Previous approaches  [58, 39] use the Sinkhorn algorithm [45] to match superpoints, which however is sensitive to initialization parameters and requires careful tuning.
+In  [46], Dual-Softmax is adopted for feature matching between two input images by using kepoints which have softmax correlation scores greater than a certain threshold.
+The correspondences of superpoints, however, are inevitably noisy due to either the noisy sensory data or simply incorrect estimations of correspondences, which prevent existing approaches from directly matching superpoints for point cloud registration.
+Postprocessing refinement is usually needed, for instance, using RANSAC-like approaches to prune the outliers [11, 12, 63, 61].
+But such refinement is inherently slow due to the iterative nature of the RANSAC pipeline and can not be easily integrated into an end-to-end trainable system.
+The recent work [59] populates another strategy, which uses a two-layer MLP to direct predict matched superpoints between the input point clouds.
+But such predicted correspondences may not exist in the point cloud, leading to inferior registration accuracy.
+In this paper, we present a strong baseline for point cloud registration by directly matching superpoints.
+Specifically, building upon RegTR [59], instead of using a MLP to predict correspondences of superpoints, we find the correspondences by computing the similarity scores of all superpoints across the source and target point cloud in a global matching manner.
+Their normlized matching scores can be used to filter out the unreliable correspondences (i.e., outliers).
+By integrating the weights of the rest inlier superpoint matchings into a differentiable variant of the Kabsh-Umeyama Algorithm [50, 23], we obtain robust estimations of the $SE(3)$ transformation between the input point cloud pairs.
+An illustration is shown in Fig.
+1.
+As a result, no ad hoc postprocessing refinement is needed, yielding a more efficient model.
+More importantly, the entire model can be trained in an end-to-end manner, where the feature representation learning, superpoints matching, and transformation estimation can be jointly optimized.
+Better registration accuracy can thus be obtained.
+In addition, compared to [59], our approach find correspondences of superpoints by global matching, which does not output non-existing points in the target point cloud.
+We run experiments on three benchmark datasets, including ModelNet [54], 3DMatch [63], and KITTI [16], and achieve comparable or even better results than state-of-the-art methods.
+We do not advocate our approach to be the solution for point cloud registration.
+Rather, we’d like to emphasize the role of matching strategy for point cloud registration, showcasing the feasibility of achieving high accuracy without cumbersome ad hoc postprocessing.
+By releasing the code and model weights, we hope our work can foster future research.
+Figure 2: Model Architecture: The KPConv backbone downsamples the input point cloud and generates superpoints and their feature vectors.
+These superpoint features are then conditioned on the other point cloud in the feature enhancement block.
+Lastly, superpoint features can be directly matched using Global Softmax to estimate SE(3) transformation using the corrleation weights in a robust manner.
+The intensity of the red and blue colors of the superpoints after the softmax correlation step represents the correspondence weights.
+(Best viewed in color.)
+II Related Work
 Traditional registration approaches.
@@
 A lot of variants [41, 9, 43, 6, 40, 30] have been proposed to improve ICP.
-Another line of work [27, 17] formulates the point cloud registration as probability distribution matching problems to improve the robustness and to converge quickly.
-However, they still heavily rely on an appropriate initialization and can easily converge to a local optimum as opposed to the global solution.
-Unlike the local methods mentioned above, global approaches are invariant to the initialization.
-Based on branch-and-bound techniques, several works [61, 8, 9, 35] have been proposed, but they are very slow and impractical in some scenarios.
-Another method is to extract and match keypoints based on feature extraction methods such as FPFH [47] and SHOT [52].
-Then RANSAC [18] can be used for registration.
-However, RANSAC is computationally very slow compared to ICP.
+Another method is to extract and match keypoints based on feature extraction methods such as FPFH [42] and SHOT [48], followed by an outlier rejection postprocessing step.
 Learning-based registration approaches.
@@
 Another line of work [8, 62] has included the transformation computation into the training pipeline.
-Following the idea of ICP [5],  [57, 56, 63, 19, 31] iteratively predict the soft correspondences and computes the transformation with differentiable weighted SVD.
-Another way [2, 33, 25, 59, 48] first extracts a global feature vector for each point cloud and predicts the transformation with global feature vectors.
-This approach usually fails in large-scale scenes.
-Unlike these works which require eighter ad-hoc postprocessing or coarse-to-fine registration, our method directly matches the superpoints without any refinement.
+Unlike these works which require either ad-hoc postprocessing or coarse-to-fine registration, our method directly matches the superpoints without any refinement.
+Learning visual correspondence.
+Various approaches have been proposed to establish correspondences between the input, e.g., using nearest neighbor followed by the distance ratio test [29].
+Recent approaches [58, 39] use the Sinkhorn algorithm [45] to match superpoints for point cloud registration.
+However, it is sensitive to initialization parameters and requires careful tuning.
+In  [46], Dual-Softmax is adopted for feature matching between two input images by using kepoints which have softmax correlation scores greater than a certain threshold.
+Global Softmax is used in a recent work of optical flow [56] to find global matchings by simply taking the correspondence for each pixel with the highest correlation score.
+In this paper, we extend Global Softmax by using the corrleation weights for both filtering unreliable correspondences and weighing the remaining inliers when estimating the SE(3) transformation matrix.
 Correspondence filters.
@@
 However, RANSAC is relatively slow and cannot be incorporated into the training pipeline because the hypothesis selection step is non-differentiable.
-To alleviate these problems, DSAC [7] modify the RANSAC pipeline and make it differentiable.
+To alleviate these problems, DSAC [7] and $\nabla$-RANSAC [53] modify the RANSAC pipeline and make it differentiable.
+But even the differentiable versions are similar in computational complexity as vanilla RANSAC.
 Other deep robust estimators [3, 10, 35, 17, 25, 60] usually use the classification network to identify which correspondences are outliers and then reject them.
 Instead of using these complex correspondence filters, our method can directly filter out outliers effectively by leveraging the rich information in the superpoints matching.
-3 Method
-Given the source and target point clouds $\mathbf{X}\in\mathbb{R}^{M\times 3}$ and $\mathbf{Y}\in\mathbb{R}^{N\times 3}$, our goal is to determine the optimal rigid transformation $\mathbf{T}=\{\mathbf{R},\mathbf{t}\}$ with rotation $\mathbf{R}\in SO(3)$ and translation $\mathbf{t}\in\mathbb{R}^{3}$ to align two point clouds into a common coordinate system.
-$M$ and $N$ denote the numbers points.
-3.1 Superpoints Feature Extraction and Enhancement
-In our approach, we use Kernel Point Convolution (KPConv) [51] as the backbone to selectively downsample the point cloud into a set of superpoints and extract global feature vectors for each superpoint.
-Specifically, the KPConv backbone uses a series of ResNet-like blocks [23] and convolutions to downsample the input point clouds into a reduced set of superpoints $\mathbf{X}{{}^{\prime}}\in\mathbb{R}^{M^{\prime}\times 3}$ and $\mathbf{Y}{{}^{\prime}}\in\mathbb{R}^{N^{\prime}\times 3}$, where $M^{\prime}<M$ and $N^{\prime}<N$.
+III Method
+Given the source and target point clouds $\mathbf{X}\in\mathbb{R}^{M\times 3}$ and $\mathbf{Y}\in\mathbb{R}^{N\times 3}$, our goal is to determine the SE(3) transformation $\mathbf{T}=\{\mathbf{R},\mathbf{t}\}$ with rotation $\mathbf{R}\in SO(3)$ and translation $\mathbf{t}\in\mathbb{R}^{3}$ to align two point clouds into a common coordinate system.
+$M$ and $N$ denote the numbers of points.
+III-A Superpoints Feature Extraction and Enhancement
+Following [59], we use Kernel Point Convolution (KPConv) [47] as the backbone to selectively downsample the point cloud into a set of superpoints and extract global feature vectors for each superpoint.
+The KPConv backbone uses a series of ResNet-like blocks [20] and convolutions to downsample the input point clouds into a reduced set of superpoints $\mathbf{X}{{}^{\prime}}\in\mathbb{R}^{M^{\prime}\times 3}$ and $\mathbf{Y}{{}^{\prime}}\in\mathbb{R}^{N^{\prime}\times 3}$, where $M^{\prime}<M$ and $N^{\prime}<N$.
 The superpoints are described by their feature vectors $\mathbf{F}_{\mathbf{X}{{}^{\prime}}}\in\mathbb{R}^{M^{\prime}\times D}$ and $\mathbf{F}_{\mathbf{Y}{{}^{\prime}}}\in\mathbb{R}^{N^{\prime}\times D}$, respectively, with $D$ being the feature dimension.
 The network weights are shared among the two point clouds.
-We use a shallower backbone for 3DMatch dataset compared to [64, 24] to avoid significant downsampling.
-Figure 2: Model Architecture: The KPConv backbone downsamples the input point cloud and generates superpoints and superpoint feature vectors.
-These superpoint features are then conditioned on the other point cloud in the Feature Enhancement block.
-Lastly, superpoint features can be directly matched using Global Softmax to predict correspondences and estimate rigid transformation.
+We use a shallower backbone for 3DMatch dataset compared to [59, 21] to avoid significant downsampling by removing the 4-th residual block.
 Although KPConv backbone provides reasonably good representations, the superpoints features are obtained within each point cloud independently.
-To obtain highly discriminative feature representations for superpoints matching, we further enhance their feature representations in the source point cloud to be conditioned on the target one and vice-versa.
-Following the previous work of point cloud registration [64] and the 2D counterpart of superpoints matching (e.g., optical flow) [60], we adopt the multi-head attention mechanism in the Transformer model [54] as the feature enhancement module, shown in Fig.
+To obtain highly discriminative feature representations for superpoints matching, We use the multi-head attention mechanism in the Transformer model [51] as the feature enhancement module as suggested in [59], shown in Fig.
 2.
@@
 The strongly associated features are strengthened while the weakly associated features are weakened.
-With such highly discriminative feature representations, we can obtain high-quality superpoints matchings.
-3.2 Direct Superpoint Matching for Rigid Transformation Estimation
+III-B Superpoint Matching for SE(3) Transformation Estimation
+Unlike [59] that uses a two-layer MLP to predict corresponding points, we find the correspondences of superpoints in a global matching manner.
+This change removes the ambiguity of predicting non-existing corresponding points.
 To get the correspondences between two point clouds, we first compare the feature similarity for each point in $\bar{\mathbf{F}}_{\mathbf{X}{{}^{\prime}}}$ to all points in $\bar{\mathbf{F}}_{\mathbf{Y}{{}^{\prime}}}$ by computing their correlations [56], which can be done efficiently in a single step as follows:
-${\mathbf{C}}=\mathrm{softmax}\left(\bar{\mathbf{F}}_{\mathbf{X}{{}^{\prime}}}\bar{\mathbf{F}}_{\mathbf{Y}{{}^{\prime}}}^{T}\right)\in\mathbb{R}^{M^{\prime}\times N^{\prime}},$ (1)
-where $\mathbf{C}$ is the normalized correlation matrix that represents the similarity between two point clouds.
-Based on the correlation matrix, the correspondences $\mathbf{\hat{Y}}$ and $\mathbf{\hat{X}}$ can be directly calculated by using the largest correlation for each point.
-The rigid transformation between the source and target point clouds can be then estimated using the superpoints correspondences with a weighted variant of the Kabsch-Umeyama algorithm [28, 53]:
+${\mathbf{C}}=\mathrm{Softmax}\left(\bar{\mathbf{F}}_{\mathbf{X}{{}^{\prime}}}\cdot\bar{\mathbf{F}}_{\mathbf{Y}{{}^{\prime}}}^{T}\right)\in\mathbb{R}^{M^{\prime}\times N^{\prime}},$ (1)
+where $\mathbf{C}$ is the normalized correlation matrix representing the similarity between two point clouds.
+Based on the correlation matrix, the correspondences between $\mathbf{X}{{}^{\prime}}$ and $\mathbf{Y}{{}^{\prime}}$ can be directly calculated by using the largest correlation for each point.
+The SE(3) transformation between the source and target point clouds can be then estimated using the superpoints correspondences with a weighted variant of the Kabsch-Umeyama algorithm [23, 50]:
 $\mathbf{\hat{R},\hat{t}}=\operatorname*{arg\,min}_{\mathbf{R},\mathbf{t}}\sum_{i}^{\min(N^{\prime},M^{\prime})}{w_{i}\lVert\mathbf{R}\mathbf{\hat{x}}_{i}+\mathbf{t}-\mathbf{\hat{y}}_{i}\rVert}^{2},$ (2)
-where $\mathbf{\hat{x}}_{i}$, $\mathbf{\hat{y}}_{i}$ are two matched superpoints, denoting the $i^{th}$ point in $\mathbf{\hat{X}}$, $\mathbf{\hat{Y}}$, respectively.
-The coefficient $w_{i}$ can be used to weigh different correspondences.
+where $w_{i}=\mathbf{C}(\hat{\mathbf{x}}_{i},\hat{\mathbf{y}}_{i})$ is the correspondence weight, $\mathbf{\hat{x}}_{i}$, $\mathbf{\hat{y}}_{i}$ are the $i$-th pair of matched superpoints.
+We use $\min(N^{\prime},M^{\prime})$ operation because corresponding points can only be found for the point cloud with the minimum number of superpoints.
 Although we use the highly discriminative feature representations enhanced by the attention module, the correspondences are inevitably noisy.
@@
 We show that the normalized correlation matrix $\mathbf{C}$ obtained from superpoints matching in Eq.
-(1) contains rich information, which allows us effectively select high-confident superpoints correspondences to estimate the transformation.
+(1) contains rich information, allowing us to effectively reject outliers for robust transformation matrix estimation.
 Specifically, if $\hat{\mathbf{x}}_{i}$ is similar to multiple superpoints, e.g., because of the repetitive patterns, its matching to $\hat{\mathbf{y}}_{i}$ tends to be unreliable.
-Therefore, the normalized correlation score between them will be low since $\mathbf{C}$ is normalized w.r.t.
+Therefore, the normalized correlation score between them $w_{i}=\mathbf{C}(\hat{\mathbf{x}}_{i},\hat{\mathbf{y}}_{i})$ will be low since $\mathbf{C}$ is normalized w.r.t.
 all other superpoints in the target point cloud.
-We can therefore set $w_{i}=\mathbf{C}(\hat{\mathbf{x}}_{i},\hat{\mathbf{y}}_{i})$, capturing the confidence of the superpoints matchings.
-On the other hand, we can use $w_{i}$ to select only highly confident superpoint matchings, e.g., top 15% of them, to estimate the transformation in Eq.(2.
-We can also augment the weights $w_{i}$ by considering the overlap score, which is introduced below, to further improve the accuracy.
-Compared to previous work [44, 66, 38], our approach is simple yet effective, which does not only eliminates the coarse-to-fine strategy but more importantly, the slow and non-differentiable RANSAC.
-As a result, our approach is fast and the superpoints feature learning, superpoints matching, and transformation estimation can be optimized jointly in an end-to-end manner, leading to superior accuracy.
+Unlike [56], which completely discards the correlation weights in $\mathbf{C}$, we use $w_{i}$ for two purposes: discarding the unreliable corresondences that have low $w_{i}$ values and further weighting the rest inliers when fitting the transformation matrix in Eq.
+(1).
+It is important to note here that $w_{i}$ is not learned and is just the correlation score between the feature representations of $\mathbf{\hat{x}}_{i}$, $\mathbf{\hat{y}}_{i}$, the $i$-th pair of matched superpoints.
+We show in the experiments that such a matching strategy works more effectively for superpoints matching than other approaches [49, 56, 46].
+In comparison to  [39, 61, 33], our approach is simple yet effective, which eliminates the coarse-to-fine strategy and more importantly, the inherently slow RANSAC-like postprocessing.
 Although the feature enhancement module is also used in [59], our approach is fundamentally different.
-Instead of predicting the correspondences by using the feature representations of the source point cloud only, we perform superpoints matching and use the rich information in the matching step to effectively filter outliers and get a robust estimation of the transformation.
-3.3 Loss Functions
+[59] predicts corresponding points for both source and target point clouds whereas our approach removes the ambiguity of predicting non-existing points and directly matches superpoints.
+The limitation of [59] can also be seen in outdoor LiDAR experiments like KITTI.
+As shown in Table III, it performs significantly worse due to predicting correspondences that do not align well with true corresponding points.
+III-C Loss Functions
 We train our approach using the following three loss functions, where the transformation loss is the main loss term and the other two are auxiliary ones.
-Transformation Loss. We apply a L1 loss on the predicted transformed locations of all keypoints with the predicted and ground truth transformation matrix.
+Transformation Loss.
+We apply the L1 loss on the predicted transformed locations of all keypoints with the predicted and ground truth transformation matrix.
+This is different from [59].
+Since we do not predict corresponding points for each point cloud, we do not need to compute two transformation matrices and thus we do not have to force the network to learn that both matrices are inverses of each other.
 $\mathcal{L}_{T}=\frac{1}{M^{\prime}}{\sum_{i}^{M^{\prime}}}\left\lvert\hat{\mathbf{R}}\mathbf{x_{i}^{\prime}}+\hat{\mathbf{t}}-\left(\mathbf{R}_{gt}\mathbf{x_{i}^{\prime}}+\mathbf{t}_{gt}\right)\right\rvert_{1}.$ (3)
@@
 Inspired by [59], we estimate the overlap values $\mathbf{\hat{O}_{\mathbf{X^{\prime}}}}$ and $\mathbf{\hat{O}_{\mathbf{Y^{\prime}}}}$ using a separate MLP layer based on the enhanced feature $\bar{\mathbf{F}}_{\mathbf{X}{{}^{\prime}}}$ and $\bar{\mathbf{F}}_{\mathbf{Y}{{}^{\prime}}}$, respectively.
-The overlap estimation is formulated as a binary classification problem, so we use binary cross-entropy loss:
-$\mathcal{L}^{X}_{o}=-\frac{1}{M^{\prime}}\sum_{i}^{M^{\prime}}{o_{x,i}^{*}\cdot\log\hat{o}_{x,i}^{*}+(1-o_{x,i}^{*})\cdot\log\left(1-\hat{o}_{x,i}\right)},$ (4)
-where $\hat{o}_{x,i}$ is the estimated overlap probability and $o_{x,i}^{*}$ is the ground truth label. We compute the overlap loss $\mathcal{L}^{Y}_{o}$ for the target point cloud similarly.
+The overlap estimation is formulated as a binary classification problem, so we use the binary cross-entropy loss:
+$\mathcal{L}^{X}_{o}=-\frac{1}{M^{\prime}}\sum_{i}^{M^{\prime}}{o_{x,i}^{*}\cdot\log\hat{o}_{x,i}+(1-o_{x,i}^{*})\cdot\log\left(1-\hat{o}_{x,i}\right)},$ (4)
+where $\hat{o}_{x,i}$ is the estimated overlap probability and $o_{x,i}^{*}$ is the ground truth probability. We compute the overlap loss $\mathcal{L}^{Y}_{o}$ for the target point cloud similarly.
 Feature Loss.
@@
 $\mathbf{W}=\mathbf{U}+\mathbf{U}^{T}$.
-The final loss is a weighted sum of all the losses above with
+The final loss is a weighted sum of all the losses with
 $\mathcal{L}=\mathcal{L}_{T}+\alpha\mathcal{L}_{f}+\beta(\mathcal{L}_{o}^{X}+\mathcal{L}_{o}^{Y}),$ (6)
 where we set the loss weights $\alpha=0.1$ and $\beta=1$ empirically.
-4 Experiments
-We evaluate our approach on two benchmarks with overlap ranging from 10% to 75%.
-The first dataset we evaluate is on ModelNet with two benchmarks settings following [24, 64] and the second dataset is 3DMatch [68] on two benchmarks following [24, 64, 44, 66].
-4.1 Implementation details
+IV Experiments
+We evaluate our approach on three datasets with overlap ranging from 10% to 75%.
+The first dataset is on synthetic ModelNet dataset with two benchmarks settings following [21, 59].
+The second one is 3DMatch [63] with two benchmarks following [21, 59, 39, 61].
+The last one is on the challenging large-scale outdoor KITTI dataset [16]
+IV-A Implementation details
 Our approach is implemented using the PyTorch framework [36] on a system with an Intel i9-1300K CPU and a single RTX 3090 GPU.
-The network training is performed with the AdamW optimizer [36], utilizing a learning rate of 0.0001 and a weight decay of 0.0001.
-For the ModelNet dataset, the network is trained for 400 epochs with a batch size of 4.
-On the 3DMatch dataset, the network is trained for 50 epochs with a batch size of 4 as well.
-Training the network requires approximately 22 hours on ModelNet and around 2 days on 3DMatch.
-4.2 ModelNet and ModelLoNet Benchmarks
-The ModelNet40 [58] dataset comprises of synthetic CAD models.
-Following the data setting in [24, 64], the point clouds are randomly sampled from mesh faces of the CAD models, cropped and subsampled.
-For the ModelNet and ModelLoNet benchmarks, the average overlap is 73.5% and 53.6%, respectively.
-Our network is trained exclusively on ModelNet, and we evaluate its generalization performance on ModelLoNet.
-For benchmarking the performance of our model we use the Relative Rotation Error (RRE) and Relative translation Error (RTE) as the primary metrics.
-Following [64], we also calculate the Chamfer Distance (CD) between the registered scan pairs.
-Table 1: Registration results on the ModelNet and ModelLoNet benchmarks.
+The network training is performed with the AdamW optimizer [31], using a learning rate of 0.0001 and a weight decay of 0.0001.
+We tain the network for 400 epochs with batch size of 4 on ModelNet, 50 epochs with batch size of 4 on 3DMatch, and 80 epochs with batch size of 1 on KITTI.
+IV-B ModelNet and ModelLoNet Benchmarks
+The ModelNet40 [54] dataset comprises of synthetic CAD models. Following the data setting in [21, 59], the point clouds are randomly sampled from mesh faces of the CAD models, cropped and subsampled.
+Our network is trained exclusively on ModelNet, and evaluated for generalization on ModelLoNet.
+For benchmarking the performance of our model we use the Relative Rotation Error (RRE) and Relative translation Error (RTE) and Chamfer Distance (CD) as the primary metrics, following [59].
+TABLE I: Registration results on ModelNet and ModelLoNet.
 Methods ModelNet ModelLoNet
 […]
-PointNetLK 29.725 0.297 0.0235 48.567 0.507 0.0367
-OMNet 2.947 0.032 0.0015 6.517 0.129 0.0074
-DCP-v2 11.975 0.171 0.0117 16.501 0.300 0.0268
-RPM-Net 1.712 0.018 0.00085 7.342 0.124 0.0050
-Predator 1.739 0.019 0.00089 5.235 0.132 0.0083
-RegTR 1.473 0.014 0.00078 3.930 0.087 0.0037
-Geotransformer 1.568 0.018 - 3.809 0.102 -
+PNLK [2] 29.725 0.297 0.0235 48.567 0.507 0.0367
+OMNet [55] 2.947 0.032 0.0015 6.517 0.129 0.0074
+DCPv2 [31] 11.975 0.171 0.0117 16.501 0.300 0.0268
+RPMNet [58] 1.712 0.018 0.00085 7.342 0.124 0.0050
+Predator [21] 1.739 0.019 0.00089 5.235 0.132 0.0083
+RegTR [59] 1.473 0.014 0.00078 3.930 0.087 0.0037
+GeoT [39] 1.568 0.018 - 3.809 0.102 -
 Ours 1.247 0.011 0.00074 3.809 0.088 0.0040
-The results are shown in Table 1.
+The results are shown in Table I.
 We compare against correspondence-based approaches [21, 59], coarse-to-fine registration approaches [39], and end-to-end methods [2, 58, 55].
-We see that our approach performs well on both benchmarks.
-Moreover, the low chamfer error suggests that predicted correspondences have very high accuracy.
-Our approach is also able to outperform methods using post-processing steps like RANSAC [24] by a significant margin, which further strengthens the point that direct superpoints matchings can actually work well for point cloud registration.
-4.3 3DMatch and 3DLoMatch Benchmarks
+Our approach performs well on both benchmarks improving results on ModelNet benchmark by 15% in RRE and 21% in RTE.
+The low chamfer error suggests that predicted correspondences have very high accuracy.
+Our approach is also able to outperform methods using post-processing steps like RANSAC [21] by a significant margin.
+IV-C 3DMatch and 3DLoMatch Benchmarks
 3DMatch [63] is a collection of 62 scenes, from which we use 46 for training, 8 for testing, and 8 for validation following [59, 21, 39].
@@
 Following the literature [21, 39, 63], We report the results of 3DMatch dataset on 5 metrics including RRE, RTE, Registration Recall (RR), Feature Matching Recall (FMR), and Inlier Ratio (IR).
-We compare our approach against several approaches including learned corresponding based algorithms [21, 13, 4, 24] and coarse-to-fine approaches [66, 44].
-Since the algorithms which use some post-preprocessing steps like RANSAC perform well only with a large number of interest points, we only show results for the maximum number of sampled points (5000).
-We also compare against several approaches designed to avoid RANSAC [10, 59, 12].
-Table 2: Registration results on the 3DMatch and 3DLoMatch benchmarks.
+TABLE II: Registration results on 3DMatch and 3DLoMatch.
 Methods 3DMatch 3DLoMatch
 […]
-3DSN 2.199 0.071 78.4 - - 3.528 0.103 33.0 - -
 FCGF 1.949 0.066 85.1 97.4 56.8 3.147 0.100 40.1 76.6 21.4
@@
 Predator 2.029 0.064 89.0 96.6 58.8 3.048 0.093 59.8 78.6 26.7
-OMNet 4.166 0.105 35.9 - - 7.299 0.151 8.4 - -
 DGR 2.103 0.067 85.3 - - 3.954 0.113 48.7 - -
-PCAM 1.808 0.059 85.5 - - 3.529 0.099 54.9 - -
 RegTR 1.567 0.049 92.0 - - 2.827 0.077 64.8 - -
@@
 CofiNet - - 89.3 98.1 49.8 - - 67.5 83.1 24.4
-Geo-Transformer (SVD) - - 86.5 - - - - 59.9 - -
-GeoTransformer (LGR) 1.625 0.053 91.5 97.7 70.3 2.547 0.074 74.0 88.1 43.3
+GeoT 1.625 0.053 91.5 97.7 70.3 2.547 0.074 74.0 88.1 43.3
 Ours 1.436 0.045 93.7 96.5 89.8 2.553 0.074 65.0 76.5 57.5
-The quantitative results are shown in Table 2.
-For the 3DMatch benchmark, our approach outperforms all the previous methods.
+We compare our approach against several learned correspondence-based algorithms [18, 11, 4, 21] and coarse-to-fine approaches [61, 39].
+The quantitative results are shown in Table II.
+For the 3DMatch benchmark, our approach outperforms all the previous methods in all but the FMR metric.
 This implies that in cases of significant overlap (>30%), the superpoint correspondences are very distinctive and accurate.
-One of the problems associated with superpoint matching in point clouds is the resolution issue, where the correspondences might not actually coincide with each other when transformed with the ground truth transformation matrix due to the subsampling of the point cloud.
-In our case, the resolution issue is automatically handled by the correlation weights.
-We also show systematic comparisons with other approaches in Table 3.
-We can see our model runs fast and is compact.
-Table 3: Comparisons of inference speed and number of parameters on the 3DMatch dataset.
-Methods Matching e2e Time Param
-CoFiNet coarse-to-fine 1.922s 5.5M
-GeoTransformer coarse-to-fine 0.088s 9.8M
-RegTR 0.063s 11.8M
-Ours direct 0.073s 7.8M
-For the 3DLoMatch benchmark, in comparison to all the approaches that do not use any post-processing, our method performs significantly better than the other approaches.
-Note that [44] uses LGR as the refinement step to get the best results, but when using all the predicted correspondences with their scores and applying weighted SVD, they get much worse results.
-This implies that the final set of all the correspondences is not optimal and contains outliers.
-In our case, these outliers are filtered by the overlap and correlation values, thus providing accurate correspondences.
-The validity of the correspondences predicted by our approach can be verified by comparing the mean Inlier Ratio in Table 2.
+For the 3DLoMatch benchmark, in comparison to all the approaches that do not use any post-processing, e.g.
+Predator[21] and CoFiNet[61], our method performs significantly better.
+GeoT [39] uses Local to Global Refinement (LGR) as the refinement step to get best results.
+In our case, outliers are filtered by the correlation weights, thus providing accurate correspondences.
+The validity of the correspondences obtained by our approach can be verified by comparing the mean Inlier Ratio (IR) in Table II.
 We get the highest mean IR, almost 19% better than the second-best approach on 3DMatch and 14% higher in 3DLoMatch.
-4.4 Ablation Studies
-Table 4: Comparisons of different outlier filtering methods.
-Outlier Filtering Method 3DMatch 3DLoMatch
-RRE RTE RR RRE RTE RR Time
-No filtering 1.462 0.046 93.3 2.652 0.074 64.6 0.073
-Ours + RANSAC-50k 2.788 0.090 82.3 4.653 0.115 32.0 0.141
-Ours (top 15%) + RANSAC-50k 1.701 0.051 93.0 2.696 0.078 64.5 0.153
-Iterative refinement 1.404 0.043 93.8 2.592 0.071 65.7 0.078
+Our approach does not perform the best in FMR because we prioritize outputting correct correspondences over a large number of, yet potentially noisy matchings.
+In contrast, CofiNet [61], has a very high FMR but low IR due to its propensity for producing many inaccurate matches.
+One of the problems associated with superpoint matching in point clouds is the resolution issue, where we might not have one-to-one correspondences due to subsampling.
+In our case, this is automatically handled by the correlation weights.
+The correlation weights are lower for weakly matching superpoints and thus even if there is no one-to-one correspondence, weak correspondences will be used according to their weights to compute the correct SE(3) transformation.
+Figure 3 shows the correlation weight distribution on the 3DMatch and 3DLoMatch benchmark.
+In 3DMatch, we have many high correlation weights, suggesting that many sample points are good correspondences.
+On 3DLoMatch, we have much fewer good correlations suggesting that many of the sampled superpoints are not good matches, which is expected because of very low overlap.
+Refer to caption Refer to caption
+Figure 3: Correlation weight distribution on 3DMatch.
+Here, many points exhibit moderate correlations, suggesting that a considerable portion of sample points are reasonably good matches.
+In contrast, 3DLoMatch has fewer points that exhibit strong correlations due to the low data overlap.
+Figure 4: Qualitative Results on 3DMatch.
+The source cloud is shown in blue with the red color representing the matched superpoints.
+Similarily target cloud is shown in yellow color with matching superpoints in blue.
+(Best viewed in color.)
+IV-D KITTI Benchmark
+KITTI odometry consists of 11 sequences of outdoor driving scenarios scanned using the HDL64 LiDAR sensor.
+Following  [21], we use sequence 0-5 for training, 6-7 for validation, and 8-10 for testing.
+As in  [21, 39] we only use the point cloud pairs that are at least 10m apart and their ground-truth poses are refined using ICP for all evaluations.
+Table III shows the results obtained on KITTI.
+We can see that the baseline [59] performs significantly worse in RTE.
+Since the correspondences are predicted, they do not align one-on-one with actual correspondences, thus significantly reducing registration accuracy.
+We achieve state of the art results in RRE and RR while performing comparably in RTE.
+This showcases the effectiveness of our approach in large scale outdoor point clouds.
+Figure 1 shows an example of matched superpoints on KITTI dataset, our approach is able to find very distinctive correspondences to register the point cloud pair.
+TABLE III: Registration results on KITTI.
+[…]
+3DFeat [57] 25.9 0.25 96.0
+FCGF [42] 9.5 0.30 96.6
+D3Feat [4] 7.2 0.30 99.8
+Predator [21] 6.8 0.27 99.8
+CofiNet [61] 8.2 0.41 99.8
+RegTR [59] 49.9 0.61 99.1
+GeoTransformer [39] 6.8 0.24 99.8
+Unified BEV [27] 7.5 0.26 99.8
+Ours 7.2 0.24 99.8
+IV-E Ablation Studies
+TABLE IV: Comparisons of different outlier filtering methods.
+Outlier Filtering 3DMatch 3DLoMatch
+Method RRE RTE RR RRE RTE RR Time
+No filtering 1.462 0.046 93.4 2.652 0.074 64.6 0.073
+Ours + RANSAC 2.788 0.090 82.3 4.653 0.115 32.0 0.141
+Ours (top 15%) + RANSAC 1.701 0.051 93.0 2.696 0.078 64.5 0.153
 Correlation scores (top 15%) 1.436 0.045 93.7 2.553 0.074 65.0 0.073
-Refer to caption Refer to caption
-Figure 3: Left: Effect of increasing the number of attention layers. Right: Comparisons of outlier filtering strategies.
-Effectiveness of the feature enhancement.
-The feature enhancement module is essential in our approach.
-To assess its effectiveness, we trained various networks with attention layers ranging from 0 to 6.
-Figure 3 illustrates the performance changes as we increase the number of attention layers.
-When no feature enhancement is used, the model’s performance is significantly poorer, which is understandable since the source point cloud’s superpoint features lack information about the target point cloud.
-Performance stabilizes at around 6 attention layers.
-It is worth noting that increasing the number of layers also leads to higher computation burden.
-Using correlation scores for filtering out outliers.
-To filter out outliers in the predicted correspondences we primarily use the correlation values along with overlap values as weights for weighted SVD.
-Since these weights are much lower for outlier values, the registration results are very accurate.
-If we use unweighted SVD (with equal weights for all the correspondences), the performance drops rapidly with the increase in number of correspondences, as shown in Fig.
-1.
-We experiment with a simple iterative refinement scheme proposed in [44], where we can iteratively re-estimate the transformation by pruning out outliers.
-We found that the performance of this refinement saturated after 5 iterations.
-See Table 4 row 4 for results.
-Fig.
-1 also shows that this approach is quite stable in pruning outliers on different number of correspondences and improves the results by a slight margin.
-Table 5: Effectiveness of different loss terms.
+TABLE V: Effectiveness of different loss terms.
 Overlap Feature 3DMatch 3DLoMatch
@@
 1.436 0.045 93.7 2.553 0.074 65.0
-We also experiment with RANSAC on the set of superpoints correspondences, See Table 4 row 2, 3.
-But as expected, RANSAC gives slightly worse results while taking more computation time.
-This implies that using our approach is much more effective in pruning outliers thus eliminating the need for refinement.
+Using correlation scores for filtering outliers.
+We study the effectiveness of outlier filtering in Table IV.
+In the ‘No filtering’ approach, we use all the superpoints with their scores without any filtering.
+A little surprisingly, it achieves very competitive results, likely because we use the correlation weights of the matchings in the transformation matrix estimation in Eq.(2).
+If the outlier filtering is used (the last row), the best accuracy can be obtained.
+We also experiment with RANSAC on the set of superpoints correspondences with and without outlier filtering, respectively.
+Without any filtering (‘Ours + RANSAC’), the large number of outliers impose significant challenges for RANSAC to find the optimal solution.
+Even with filtering (‘Ours (top 15%) + RANSAC’), we can see that RANSAC gives slightly worse results, partially because it does not always converge to the optimal solution.
+Furthermore, RANSAC requires twice as much time for the estimation of the transformation matrix compared to our method.
 Effectiveness of the loss terms.
-Lastly, we analyze the effectiveness of each loss function.
-Table 5 shows the results with different loss function configurations.
-We see that by just using Feature Loss with Transformation loss, the model is able to achieves good performance and using Overlap loss is the final step that helps the model prune remaining outliers.
-Row 3 in Table 5 again implies that superpoint correspondences are very distinctive.
-Figure 4: Qualitative Results on 3Dmatch dataset: The source cloud is shown in Green with the red color representing the matched superpoints.
-Similarily target cloud is in red color with matching superpoints in green color.
-5 Limitations and Broader Impacts
-We have identified two limitations in the current architecture, First, our approach requires a decent amount of superpoints in the low overlapping region to perform well.
-Second, our approach samples about 200 best correspondences, but for the low overlapping region only a few of them lie on the overlapping region thus making it hard for the network to achieve good performance on 3DLoMatch benchmark.
-Instead of regularly downsampling, structure-aware sampling strategies are needed.
-This research has the potential to significantly impact the fields of 3D computer vision and robotics, as well as any domains that rely on accurate and efficient point cloud registration, such as autonomous navigation, 3D mapping, and augmented reality.
-By introducing a simple yet effective approach for directly matching superpoints in an end-to-end manner, the research not only speeds up the registration process but also enhances its robustness by effectively filtering out outliers without any postprocessing refinement.
-6 Conclusion
-We introduce a straightforward yet highly effective method for point cloud registration that offers simplicity in understanding and implementation.
-Our approach directly matches superpoints to establish correspondences, enabling the computation of rigid transformations and correlation weights.
-To enhance performance in low overlap regions, we augment the correlation weights with predicted overlap values.
-Experimental evaluations on the 3DMatch and ModelNet datasets demonstrate the efficacy of our approach across different overlap scenarios.
-Notably, our approach is computationally efficient, robust to outliers, and does not require post-processing steps, making it suitable for real-time applications.
-In the future, we plan to extend its application to cross-modalities such as image-to-image and image-to-point cloud registration.
-Appendix A Appendix
-A.1 Network Architecture
-Figure 5: Network Architecture of the KPConv backbone for 3DMatch and ModelNet dataset.
-We do not downsample the point cloud by a large factor to maintain resolution on the overlapping region of the point cloud pair.
-Figure 6: Network architecture of the Feature Enhancement Module.
-The architecture is shown for just one block of self-attention and cross-attention layers but can be extended similarly for improvement in performance.
-The network architecture for the KPConv backbone used is shown in Fig. 5 and the architecture for the Feature Enhancement module is shown in Fig. 6.
-We do not downsample the point cloud by a large factor in the bakbone unlike other methods [44, 64] as this leads to a very low resolution in the overlapping region of the point cloud pair.
-Since, we are training the network to learn correlations between the downsampled superpoints, it is beneficial to have a moderately larger number of points.
-A.2 Evaluation Metrics
-We report Registration Recall (RR), Relative Rotation Error (RRE), Relative Translation Error (RTE), Feature Matching Recall (FMR), and Inlier Ratio (IR) for the 3DMatch dataset following common practice [44, 64, 24, 66, 4].
-For ModelNet dataset, we report RRE, RTE, and Chamfer Distance Error (CD) following [64, 24, 59].
-A.2.1 3DMatch/3DLoMatch
-Inlier Ratio is the fraction of inlier correspondences in all predicted correspondences.
-Corresponding points are considered inliers if the distance between them under the ground-truth transformation is smaller than $\tau_{1}=10cm$
+We also analyze the effectiveness of each loss function.
+Table V shows the results with different loss function configurations.
+We see that by just using the feature loss along with the transformation loss, the model is able to achieve good performance.
+Using the overlap loss further helps the model prune remaining outliers.
+Matching strategy.
+The choice of a matching strategy is crucial for precise pose estimation.
+In prior approaches, such as [39, 58, 61], the Sinkhorn algorithm has commonly been employed to compute optimal transport and derive matching scores for correspondences.
+In our research, we opt for Global Softmax, which, as evidenced by our experiments, outperforms the Sinkhorn approach (refer to Table VI).
+Note that while Sinkhorn algorithm requires careful parameter tuning, our approach is parameter-free and works out of the box.
+While Global Softmax has been previously used as a matching strategy in [49, 58], prior works only utilized it to identify matching points with probabilities exceeding a certain threshold, resulting in suboptimal outcomes.
+Our innovation lies in utilizing these probabilities as correspondence weights, facilitating gradient flow through the Weighted Kabsch-Umeyama Solver which leads to faster optimization and better results.
+We also experimented with Dual Softmax matching used in LofTR [46], but the difference in performance is less than 1% with Global Softmax being slightly better and faster than Dual-Softmax.
+TABLE VI: Comparison of Matching Strategy
+Methods 3DMatch 3DLoMatch Time
 […]
+Sinkhorn 1.509 0.046 89.2 2.622 0.074 59.6 0.094
+Dual Softmax 1.441 0.045 93.6 2.589 0.074 65.1 0.080
+MLP [59] 1.567 0.049 92.0 2.827 0.077 64.8 0.078
+Global Softmax (Ours) 1.436 0.045 93.7 2.553 0.074 65.0 0.073
+Data Augmentation Schemes Previous works have tried different data augmentation schemes.
+[59] uses a weak data augmentation, perturbing poses by a small amount during training.
+While  [39, 58, 21] use large data augmentation, perturbing poses by a full range of motion.
+Weak data augmentation can slightly improve performance but deteriorates network generalization.
+Table VII shows that our approach achieves slightly better performance with weak data augmentation.
+But with strong data augmentation, the generalization on 3DLoMatch benchmarks improves RR by 1.8%.
+TABLE VII: Comparison of Data Augmentation Schemes
+Methods 3DMatch 3DLoMatch
 […]
-Feature Matching Recall is the fraction of point cloud pairs whose inlier ratio is above $\tau_{2}=0.05$.
-It indicates the likelihood of recovering the optimal transformation between two point clouds.
-It is a good indicator for methods which use post-processing refinement steps like RANSAC [66, 44, 24] as having a higher FMR increases the probability of recovering the correct transformation.
-If the total number of point cloud pairs in the test set is $\mathcal{N}$, FMR can be calculated as
-$\textnormal{FMR}=\frac{1}{\mathcal{N}}\sum^{\mathcal{N}}_{i=1}\llbracket\textnormal{IR}>\tau_{2}\rrbracket.$ (8)
-Registration Recall represents the fraction of point cloud pairs whose Root Mean Square Error (RMSE) is less than $\tau_{3}=20cm$. RR directly measures the effectiveness of the registration algorithm.
-$\textnormal{RR}=\frac{1}{\mathcal{N}}\sum^{\mathcal{N}}_{i=1}\left\llbracket\sqrt{\frac{1}{|\mathcal{C^{*}}|}\sum_{(\mathbf{x}_{i},\mathbf{y}_{i})\in\mathcal{C^{*}}}\lVert\mathbf{T}^{\mathbf{X}}_{\mathbf{Y}}(\mathbf{x}_{i})-\mathbf{y}_{i}\rVert^{2}_{2}}<\tau_{3}\right\rrbracket,$ (9)
-where $\mathcal{C^{*}}$ is the set of ground truth correspondences and $\mathbf{T}^{\mathbf{X}}_{\mathbf{Y}}$ is the predicted transformation matrix.
-Relative Rotation Error and Relative Translation Error are the rotation and translation errors between the ground truth transformation matrix $\mkern 1.5mu\overline{\mkern-1.5mu\mathbf{T}\mkern-1.5mu}\mkern 1.5mu^{\mathbf{X}}_{\mathbf{Y}}\in SE(3)$ and the predicted transformation $\mathbf{T}^{\mathbf{X}}_{\mathbf{Y}}\in SE(3)$.
-We can obtain $\mathbf{R}\in SO(3)$, $\mkern 1.5mu\overline{\mkern-1.5mu\mathbf{R}\mkern-1.5mu}\mkern 1.5mu\in SO(3)$, $\mathbf{t}\in\mathbf{R}^{3}$ and $\mkern 1.5mu\overline{\mkern-1.5mu\mathbf{t}\mkern-1.5mu}\mkern 1.5mu\in\mathbf{R}^{3}$ as the predicted rotation matrix, ground truth rotation matrix, predicted translation vector and ground truth translation vector respectively.
-Thus, RRE and RTE can be calculated as follows
-$\textnormal{RRE}=\textnormal{arccos}\left(\frac{\textnormal{trace}(\mathbf{R}^{T}\mkern 1.5mu\overline{\mkern-1.5mu\mathbf{R}\mkern-1.5mu}\mkern 1.5mu)-1}{2}\right),~{}~{}~{}\textnormal{RTE}=\lVert\mathbf{t}-\mkern 1.5mu\overline{\mkern-1.5mu\mathbf{t}\mkern-1.5mu}\mkern 1.5mu\rVert.$ (10)
-A.2.2 ModelNet/ModelLoNet
-Chamfer Distance measures the dissimilarity or discrepancy between two sets of points or point clouds.
-It quantifies the minimum average distance between each point in one set and its nearest neighbor in the other set.
-It provides a measure of how well two point clouds align or match with each other, with lower values indicating a better alignment.
-$\text{{CD}}(\mathbf{X},\mathbf{Y})=\sum_{x\in\mathbf{X}}\min_{y\in\mathbf{Y}}\|y-\mathbf{T}^{\mathbf{X}}_{\mathbf{Y}}(x)\|_{2}^{2}+\sum_{y\in\mathbf{Y}}\min_{x\in\mathbf{X}}\|y-\mathbf{T}^{\mathbf{X}}_{\mathbf{Y}}(x)\|_{2}^{2},$ (11)
-where $\mathbf{X}$ and $\mathbf{Y}$ represent the sets of points in two point clouds, $\|\cdot\|_{2}$ denotes the Euclidean distance and $\mathbf{T}^{\mathbf{X}}_{\mathbf{Y}}$ is the predicted transformation between point clouds $\mathbf{X}$, and $\mathbf{Y}$.
-A.3 Additional Results
-Table 6 shows the scene-wise registration results on the 3DMatch and 3DLoMatch benchmarks.
-Following literature [44, 64, 24], we only calculate mean RRE and RTE for the successfully registered scan pairs.
-For the 3DMatch benchmark, our method outperforms other methods in every metric.
-Our methods achieves not only the highest registration recall, but also the lowest mean RRE and RTE.
-This implies that the for all the successfully registered points clouds, the predicted correspondences and their correlation values are very accurate.
-For the 3DLoMatch benchmark, our method outperforms every method which does not use post-processing refinement like RegTR, while giving the lowest RRE and RTE values.
-In comparison to approaches which use post-processing refinement like GeoTransformer, Predator, and CoFiNet, the performance gap is expected as the predictions in 3DLoMatch contain a much higher number of outliers.
-But the gap is minor and our approach runs faster.
-Table 6: Scene-wise registration results on the 3DMatch and 3DLoMatch benchmarks.
-Method 3DMatch 3DLoMatch
-Kitchen Home 1 Home 2 Hotel 1 Hotel 2 Hotel 3 Study Lab Mean Kitchen Home 1 Home 2 Hotel 1 Hotel 2 Hotel 3 Study Lab Mean
-Registration Recall (%)normal-$\uparrow$
-3DSN 90.6 90.6 65.4 89.6 82.1 80.8 68.4 60.0 78.4 51.4 25.9 44.1 41.1 30.7 36.6 14.0 20.3 33.0
-FCGF 98.0 94.3 68.6 96.7 91.0 84.6 76.1 71.1 85.1 60.8 42.2 53.6 53.1 38.0 26.8 16.1 30.4 40.1
-D3Feat 96.0 86.8 67.3 90.7 88.5 80.8 78.2 64.4 81.6 49.7 37.2 47.3 47.8 36.5 31.7 15.7 31.9 37.2
-Predator 97.6 97.2 74.8 98.9 96.2 88.5 85.9 73.3 89.0 71.5 58.2 60.8 77.5 64.2 61.0 45.8 39.1 59.8
-CoFiNet 96.4 99.1 73.6 95.6 91.0 84.6 89.7 84.4 89.3 76.7 66.7 64.0 81.3 65.0 63.4 53.4 69.6 67.5
-RegTR 97.8 90.6 75.5 97.8 94.9 100 88.5 91.1 92.0 66.2 58.5 64.9 72.7 61.3 70.7 53.0 71.0 64.8
-GeoT 98.9 97.2 81.1 98.9 89.7 88.5 88.9 88.9 91.5 85.9 73.5 72.5 89.5 73.2 66.7 55.3 75.7 74.0
-Ours 97.6 88.7 87.4 97.8 93.6 100 87.2 97.8 93.7 62.3 55.3 69.4 72.2 62.0 45.6 46.2 76.8 65.0
-Relative Rotation Error (°)normal-$\downarrow$
-3DSN 1.926 1.843 2.324 2.041 1.952 2.908 2.296 2.301 2.199 3.020 3.898 3.427 3.196 3.217 3.328 4.325 3.814 3.528
-FCGF 1.767 1.849 2.210 1.867 1.667 2.417 2.024 1.792 1.949 2.904 3.229 3.277 2.768 2.801 2.822 3.372 4.006 3.147
-D3Feat 2.016 2.029 2.425 1.990 1.967 2.400 2.346 2.115 2.161 3.226 3.492 3.373 3.330 3.165 2.972 3.708 3.619 3.361
-Predator 1.861 1.806 2.473 2.045 1.600 2.458 2.067 1.926 2.029 3.079 2.637 3.220 2.694 2.907 3.390 3.046 3.412 3.048
-CoFiNet 1.910 1.835 2.316 1.767 1.753 1.639 2.527 2.345 2.011 3.213 3.119 3.711 2.842 2.897 3.194 4.126 3.138 3.280
-RegTR 1.729 1.34 1.79 1.639 1.28 1.810 1.57 1.35 1.56 3.366 2.446 3.244 2.732 2.439 2.919 3.044 2.428 2.827
-GeoT 1.79 1.353 1.79 1.52 1.328 1.57 1.952 1.678 1.625 2.35 2.30 2.541 2.455 2.49 2.504 3.01 2.716 2.547
-Ours 1.780 1.272 1.596 1.457 1.239 1.456 1.492 1.196 1.436 2.330 2.281 2.59 2.72 2.520 2.58 2.775 2.61 2.55
-Relative Translation Error (m)normal-$\downarrow$
-3DSN 0.059 0.070 0.079 0.065 0.074 0.062 0.093 0.065 0.071 0.082 0.098 0.096 0.101 0.080 0.089 0.158 0.120 0.103
-FCGF 0.053 0.056 0.071 0.062 0.061 0.055 0.082 0.090 0.066 0.084 0.097 0.076 0.101 0.084 0.077 0.144 0.140 0.100
-D3Feat 0.055 0.065 0.080 0.064 0.078 0.049 0.083 0.064 0.067 0.088 0.101 0.086 0.099 0.092 0.075 0.146 0.135 0.103
-Predator 0.048 0.055 0.070 0.073 0.060 0.065 0.080 0.063 0.064 0.081 0.080 0.084 0.099 0.096 0.077 0.101 0.130 0.093
-CoFiNet 0.047 0.059 0.063 0.063 0.058 0.044 0.087 0.075 0.062 0.080 0.078 0.078 0.099 0.086 0.077 0.131 0.123 0.094
-RegTR 0.040 0.041 0.057 0.057 0.042 0.039 0.054 0.058 0.049 0.079 0.064 0.078 0.094 0.074 0.060 0.093 0.077 0.077
-GeoT 0.042 0.046 0.059 0.055 0.046 0.050 0.073 0.053 0.053 0.062 0.070 0.071 0.080 0.075 0.049 0.107 0.083 0.074
-Ours 0.043 0.037 0.061 0.051 0.044 0.029 0.048 0.044 0.045 0.061 0.068 0.068 0.088 0.071 0.054 0.091 0.091 0.074
-A.4 Iterative Pose Refinement
-We experiment with a simple iterative refinement scheme proposed in [44], where we can iteratively re-estimate the transformation by pruning out outliers.
-$\mathbf{R,t}=\max_{\mathbf{R_{i}},\mathbf{t_{i}}}\sum_{i}{\llbracket\lVert\mathbf{R}\mathbf{x}_{i}+\mathbf{t}-\mathbf{y}_{i}\rVert^{2}<\tau_{a}\rrbracket},$ (12)
-[…]
-Refer to caption Refer to caption
-Figure 7: Improvement in Reigstration Recall on increasing the number of iterations for pose refinement.
-We see that the performance saturates after the 5th iteration for 3DLoMatch and after 2 iterations for 3DMatch.
-A.5 Qualitative Results
-We provide more visual results in Fig. 8 and Fig. 9.
-Figure 8: Qualitative Results on the ModelLoNet dataset.
-Row (a)-(d) show the cases with successful registration.
-While the rows (e) and (f) show the failure cases.
-The source cloud is shown in Green with the red color representing for the matched superpoints.
-Similarily target cloud is in red color with matching superpoints in green color.
-We see that our approach only fails in cases which are very ambigous.
-Figure 9: Qualitative Results on the 3DLoMatch dataset.
-Row (a)-(d) show the cases with successful registration.
-While the rows (e) and (f) show the failure cases.
-The source cloud is shown in Green with the red color representing for the matched superpoints.
-Similarily target cloud is in red color with matching superpoints in green color.
+Ours (Weak Augmentation) 1.436 0.045 93.7 2.553 0.074 65.0
+Ours (Large Augmentation) 1.472 0.045 93.4 2.611 0.074 66.8
+V Conclusion
+In this paper, we presented a strong baseline approach for point cloud registration.
+By using Global Softmax to directly match superpoint features, we remove the ambiguity of predicting non-existent corresponding points while using the softmax probabilities as correspondence weights allows us to filter outliers without any post-processing refinement.
+Experimental results on standard benchmarks show that our model achieves comparable or even better accuracy than state-of-the-art methods.
+We do not advocate our approach to be the solution for point cloud registration.
+Rather, we’d like to emphasize the role of matching strategy for point cloud registration, showcasing the feasibility of achieving high accuracy without cumbersome ad hoc postprocessing.